A Complete Bibliography of Publications in Biometrics: 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

18 September 2024 Version 2.00

Title word cross-reference

2^{3} [504]. 2 × 2 [268, 267]. 3 × 6 [368]. 4^{n} [369]. 5 × 5 [88]. χ2 [408]. χ^{2} [266, 309, 264, 265, 103]. F [279]. F_{1} × F_{2} [407]. k [457]. k_{4} [386]. r × c [103]. × [203]. X^{2} [159].

-Visit [457].

100 [221]. 113 [287]. 1954 [304].

50 [64, 105, 549, 374, 151].

8 [578].

Abscissa [535]. abstract [573]. Abstracts [26, 81, 107, 115, 129, 138, 153, 162, 190, 236, 251, 274, 334, 288, 303, 327, 350, 364, 393, 423, 434, 450, 464, 477, 490, 506, 522, 538, 554, 572]. Accident [113, 458]. Accident-Proneness [113]. Acreage [486]. Action [294, 566]. Address [333, 216]. Adequate [548]. Adjusted [399, 428]. Adjustment [347, 340]. Advances [388]. Affinity [406]. After [501]. After-History [501]. Against [56]. Age [62, 358]. Age-Specific [358]. Agricultural [345, 474]. Allelic [344]. Allergy [239]. Alternative [361, 319]. Alternatives [253]. among [72, 91]. Analyses [430, 386]. Analysis [565, 200, 361, 448, 570, 571, 229, 564, 295, 383, 60, 257, 456, 29, 211, 323, 104, 158, 427, 433, 71, 514, 321, 119, 372, 18, 345, 381, 231, 232, 515, 214, 500, 355, 243, 77, 258, 403, 320, 419, 517, 444, 168, 445, 308, 205, 135, 180, 382, 8, 319, 513, 51, 291, 547, 55, 201, 74, 428, 462, 318, 296, 178, 530, 248, 37, 431, 484, 475, 254, 246, 429, 368, 497]. Angle [549]. Angular [211]. Animal [36, 485, 359]. Announcements [379, 394, 411, 425, 436, 452, 466, 479, 492, 508, 524, 540, 556, 575]. Answer [287]. Antigenicity [283]. any [158]. Apparatus [298]. Application [446, 100, 197, 405, 400, 38, 19, 90, 502, 357, 8, 61]. Applications [31, 442, 372, 386]. Applied [266, 461]. Appraisals [3]. Appreciation [449, 421]. Approach [564, 299]. Appropriate [322, 417]. Appropriateness [257]. Approximation [562]. Area [532, 474]. Area-Sample [532]. Arranged [321, 368]. Arrangements [119, 422]. Asay [52]. Aspect [143]. Aspects [52, 111]. Assay [6, 564, 66, 149, 197, 1, 51, 414]. Assays [295, 158, 341, 92, 65, 322, 50]. Associated [486]. Asymptotic [375, 529, 99]. Attack [438]. Attempt [283]. Australian [418]. Auto [512]. Auto-Tetraploids [512]. Automatic [159]. Average [549, 123]. Average-Angle [549]. Axis [311].

Bacteria [85, 470]. Bacterial [17, 40, 298]. Balance [432, 398]. Balanced [101, 228, 419, 86, 135, 125, 246]. Baled [76]. Bartlett [504]. Based [111]. Basic [307]. BCG [239]. Be [347]. Beetles [446, 405, 178]. Best [487]. between [122, 371, 54, 307, 553, 253, 471, 346, 255, 521]. beyond [372]. Binomial [182, 231, 232, 420, 461, 183]. Bio [6, 66, 1, 322]. Bio-Assay [6, 66, 1]. Bio-Assays [322]. Bioassay [402, 496]. Bioassays [391]. Biologic [148]. Biological [52, 564, 182, 197, 92, 50, 414]. Biologie [90]. Biology [297, 90, 194]. Biometric [33, 191, 481, 32, 542, 62, 27, 46, 68, 82, 108, 130, 163, 176, 192, 222, 237, 275, 289, 335, 351, 365, 395, 410, 424, 435, 451, 465, 478, 491, 507, 523, 539, 555, 574]. Biometrical [52, 31, 143]. Biometrics [578, 326]. Biometry [13, 12, 297, 10, 388, 352]. Birth [133]. Bliss [421, 134]. Block [101, 122, 257, 228, 104, 339, 372, 270, 258, 419, 245, 86, 135, 531, 125, 398, 317, 179, 246, 429]. Blocks [200, 500, 166, 180, 382]. Blood [89, 61, 441, 97]. Borer [404, 489]. Both [387]. Bradley [310]. Branch [285, 413]. Brief [333].

Calculating [159]. Calculation [518, 151, 103]. Calculations [374]. Calculator [136, 233]. cas [102]. case [102]. Catch [460]. Categories [417].

Causal [530]. Cell [124, 61]. Cells [441]. Censuses [156, 578]. Center [520]. Central [520]. Centric [532]. certain [102]. certains [102]. Chain [245, 179]. Chains [266]. Change [326, 519]. Change-Over [519]. Chapman [578]. Characterization [283]. Characters [91]. Charles [93]. Chemical [33, 32, 226]. Chester [421]. Chi [504, 560, 518, 551]. Chi-Square [518, 551]. Chi-Squares [504]. Chickens [201]. Children [62]. Choice [6, 1]. Chromosome [406]. Class [473, 536, 320, 98, 531]. Classes [362, 74]. Classification [321, 320, 74]. Classifications [430]. Clearance [91]. Climates [390]. Cline [59]. Clinical [149, 546, 142]. Cloud [311]. Cochran [11]. Coefficient [360, 312, 552]. Coefficients [476, 521]. Collaborative [51]. Combination [341, 230]. Combining [561]. Comments [105, 550]. Common [487, 500, 264]. Comparative [197, 281]. Comparing [195, 516, 121]. Comparison [571, 442, 208, 134, 502, 544, 403, 455, 547, 563]. Comparisons [268, 383, 257, 443, 280]. Complete [500]. Complex [381]. Component [71, 386]. Components [20, 76, 75, 186, 77, 459, 420, 292, 73]. Composite [520, 567]. Compound [37]. Computation [136]. Computers [355]. Computing [233, 444, 445]. Concentration [97]. Concept [113, 360]. Concepts [120]. Conditions [339]. Conference [191]. Confidence [391, 244, 354, 549, 535, 343, 529]. Connected [440]. Consequent [340]. Considerations [227, 148, 269]. Consistency [75]. Constant [298]. Constraints [381]. Constructing [526]. Construction [35, 123]. Consumer [565, 457]. Contagious [202, 293, 473]. Containing [254]. Content [89]. Contingency [565, 309, 504, 265, 568, 159, 342, 103]. Continuous [282, 497, 344, 483, 385]. Contrasts [534]. Contributed [57]. Contribution [550, 55, 55]. Contributions [280, 343]. Control [304, 361, 247, 553, 389, 319, 563]. Convenient [151]. Corn [404, 489]. Correction [6, 24, 199, 378, 221]. Corrections [446, 447, 448, 505, 570, 571]. Correlated [228, 416, 399]. Correlation [169, 255, 312, 552, 521]. Corrigenda [101]. Corticosteroids [62]. Cotton [486]. Count [241, 569, 285]. Counts [124, 503, 40, 61]. Course [297, 513]. Covariance [304, 427, 433, 247, 430, 432, 186, 559, 319, 428, 462, 340, 431, 429, 361]. Covariate [498]. Critique [64, 105]. Crop [471, 299, 296, 254]. Crop-Weather [299]. Crops [513, 376]. Cross [208, 502]. Cross-Comparison [502]. Cross-Over [208]. Culling [527]. Curve [356, 375, 358]. Curves [60, 29, 184, 56, 455]. Cut [309, 265].

D [578]. D’Angelo [111]. dans [216, 102]. Data [505, 564, 182, 560, 171, 119, 38, 85, 469, 77, 403, 286, 240, 296, 111, 431, 484, 475, 97]. Death [322]. Death-Times [322]. Decade [481]. Decision [487, 253]. Degrees [445]. Densities [17]. Dependent [417]. Derivation [92, 255, 484]. Derive [283]. Description [282, 483]. Design [49, 197, 520, 294, 270, 148, 306, 269, 168, 308, 205, 180, 376, 226, 528, 318, 317, 248, 179, 200]. Designs [101, 122, 487, 228, 208, 47, 35, 372, 7, 515, 567, 419, 245, 86, 135, 519, 531, 125, 398, 246, 429]. Desk [136, 233]. d’estimation [102]. Detection [458, 111]. Deterioration [438]. Determination [339, 241, 311]. Determine [148, 548]. Determined [59]. Determining [299, 285, 21, 291]. Development [546, 239]. Deviate [560]. Deviations [4]. Diagnostic [63, 98]. Diallel [243]. Difference [371, 553, 271, 281, 521]. Differences [122, 458]. Different [230, 362, 254]. Differential [537]. Diffusion [59]. Digitalis [51]. Dignités [216]. dignities [216]. Dilution [510]. d’Incubation [55]. Diphtheria [283]. Discrimination [385]. Discussion [52, 57, 49, 33, 11, 13, 31, 144, 147, 150, 232]. Distribution [182, 293, 314, 536, 255, 404, 489, 212, 370, 272, 529, 461, 441]. Distribution-Free [529]. Distributions [202, 213, 440, 473, 260, 120, 157]. Divisible [419]. Dose [198, 151]. Dose-Response [198]. Double [321, 102]. Doubly [228]. Dried [357]. Drift [344]. Drosophila [472]. Drug [97]. Drugs [199, 346, 566, 146]. Dual [125]. Due [438]. d’un [55, 311]. Dynamics [36].

échantillonnage [102]. Ecology [485]. ED [374, 151]. Editor [324]. Editorship [326]. Education [13, 12]. Effect [383, 89, 468]. Effective [151]. Effects [228, 339, 545, 390]. Efficiency [209]. Efficient [197, 471, 183]. Egg [201]. Elements [61, 132]. Elimination [245]. Engineering [18]. Enumeration [132]. Environment [497, 483]. Environmental [527]. Eosinophil [241]. Epidemic [266]. Equation [198]. Equations [100, 299, 284, 463]. Equilibria [512]. Equilibrium [54]. Errata [200, 304, 489]. Erratum [309, 361, 578]. Error [66, 149, 241, 5, 443, 387, 414, 317]. Errors [124, 486, 92, 389, 484]. Erythrocyte [38, 89]. Estimate [401]. Estimated [209]. Estimates [64, 341, 230, 75, 134, 486, 550, 459, 414, 409, 105]. Estimating [56, 413, 529, 460, 78, 359]. Estimation [505, 149, 17, 30, 5, 469, 471, 186, 527, 536, 472, 172, 87, 362, 343, 240, 120, 157, 272, 358, 475, 102]. Estimators [561]. European [404, 489]. Evaluation [167, 63, 403, 357, 98, 528, 359]. Exact [309, 265, 344]. Examination [85, 516]. Example [307, 208, 314, 196]. Examples [227, 419]. Excretion [62]. Existing [372]. Expectation [4]. Expectations [291]. Experiences [474]. Experiment [197, 514, 381, 513, 548, 368]. Experimental [49, 446, 487, 145, 405, 440, 47, 498, 306, 384, 376, 528, 28]. Experimentation [140, 48]. Experiments [304, 570, 571, 510, 442, 228, 230, 498, 247, 345, 214, 500, 253, 269, 166, 181, 422, 517, 308, 226, 547, 318, 340, 248, 37, 254, 369]. Exploitation [227, 307]. Exploration [227, 307]. Extended [208]. Extension [373, 399]. Extensions [489, 367, 447]. Extra [519]. Extra-Period [519]. Extreme [560].

Factor [488, 518]. Factorial [268, 35, 119, 55, 369]. Factorials [419, 573]. Factorielle [55]. Factors [239]. Family [239, 543, 439]. Fertilized [384]. Fertilizers [214]. Fiducial [20]. Field [404, 489, 248]. Fields [486]. First [481]. Fish [460]. Fisher [213]. Fit [551]. Fitted [307, 535]. Fitting [182, 293, 356, 495, 387, 184, 375, 183]. Fixed [229, 448]. Fixed-Sample-Size [229, 448]. Flavor [383, 3, 166, 281]. Flour [446, 405]. Fluid [241]. Fly [195, 468]. Fly-Repellant [195]. Following [33, 11, 13]. Foodstuffs [3]. Formulas [309, 265]. Foundations [485]. Four [472, 513, 39]. Four-Course [513]. Fractional [35, 196, 378, 338]. Free [56, 529]. Freedom [445]. French [216, 90, 102, 499, 55, 311]. Frequencies [313, 59, 472, 420, 342]. Frequency [61]. Front [16, 34, 58, 70, 84, 96, 109, 118, 131, 139, 154, 165, 177, 193, 207, 225, 238, 252, 263, 277, 278, 290, 305, 315, 337, 353, 366, 380, 397, 412, 426, 437, 454, 467, 480, 494, 525, 541, 557, 577]. Fruit [269, 285, 413, 21]. Function [401, 548]. Functions [213]. Fundamental [120]. Fungal [438]. Further [280, 239, 286]. Future [542].

G [11, 578]. Gene [59]. General [227, 7, 355, 120]. Generalization [202]. Generalized [473]. Genes [85]. Genetic [354, 527, 344, 292, 559, 312, 528, 552, 386, 548]. Genetics [31, 360, 284]. Genic [282]. Genotype [497, 483]. Goodness [551]. Goodness-of-Fit [551]. Gordon [111]. Graded [92, 346]. Gradients [304, 361, 247, 319]. Graduate [297]. Graphical [92]. Grasshoppers [418]. Greenhouse [422]. Gross [87]. Group [373, 419, 399, 558, 543, 439]. Grouped [550, 496]. Growth [570, 60, 185, 517, 455, 298, 256].

Haldane [499]. Half [181, 543, 548]. Half-Plaid [181]. Half-Sib [543]. Half-Sibs [548]. Hand [56]. Having [246]. Hbn. [404, 489]. Height [62]. Hemacytometer [503, 441]. Hemoglobin [89]. Heritability [354, 172, 529]. Heritable [407]. Heterogeneity [560, 270, 245]. Heteroscedastic [416]. History [501]. Hit [356, 184]. Hormones [294, 111]. Host [510]. Hot [390]. Human [56, 111]. Hypothesis [518].

II [571, 497, 157, 547]. III [272]. Illustrated [135]. Illustration [562]. Impact [360]. Improvement [28]. Inadequate [347]. Incomplete [101, 505, 257, 228, 372, 469, 166, 258, 419, 86, 135, 531, 125, 398, 317, 431, 484, 475, 246, 429]. Increase [244]. Incubation [55]. Independent [526, 521]. India [460]. Indices [511]. Individual [498, 458, 132]. Industrial [35]. Industry [33, 32, 306]. Infection [358]. Infectivity [415]. Information [69, 117, 164, 224, 276, 336, 396, 453, 509, 576, 104, 398]. Inherently [415]. Insecticide [318, 248]. Instructions [151]. Instruments [78]. Insulin [208]. Intensities [502]. Intensity [281]. Inter [104, 398, 317]. Inter- [317]. Inter-Block [104, 398]. Interaction [497, 518, 483, 203]. Interactions [54, 282, 385]. Interactive [566]. International [191, 51]. Interpolation [123]. Interpretation [428, 408]. Intersection [535]. Interval [244, 535, 343]. Intervals [20, 476, 354]. Intra [419, 317]. Intra-Block [419, 317]. Introduction [9, 330]. Introductory [297]. Invasion [198]. Inverse [156, 40, 578]. Inversely [209]. Inverted [299]. Investigation [209]. Investigations [534]. Involving [381, 513, 548].

Irrigation [181]. Item [533]. Iterative [463]. IV [550].

Japan [388]. Joint [294, 500, 255]. June [304].

Kronecker [400].

Laboratory [18, 166]. Lady [178]. l’Analyse [55]. Lancaster [504]. Larvae [404, 489]. Latin [368, 39, 316]. Latinized [126]. Lattice [7, 515, 205]. Lattices [101, 104, 5, 126, 86, 135, 173]. Laurent [493]. Law [214]. l’Axe [311]. LD [64, 105, 549, 374, 151]. Learning [545]. Lee [213]. Lesions [239]. Letter [324]. Likelihood [505, 100, 401, 469, 495, 347, 461, 463]. Limits [391, 549, 529]. Line [387, 169, 414, 485]. Linear [313, 72, 535, 420]. Link [307]. Linkage [385]. Linked [200, 180, 382]. Litchfield [134]. Little [573]. Local [239]. Locule [21]. Locus [344]. Locusts [418]. Logistic [401, 495]. Logits [112]. Low [415].

Machine [159]. macronuclear [256]. Males [468]. Many [253, 242]. Mapping [406]. Marrow [132]. Matched [371]. Materials [438]. Mathematical [54, 90, 283, 53, 485]. Mathématique [90]. Matrices [323, 400, 255]. Matrix [299]. Matter [16, 34, 58, 70, 84, 96, 109, 118, 131, 139, 154, 165, 177, 193, 207, 225, 238, 252, 263, 277, 278, 290, 305, 315, 337, 353, 366, 380, 397, 412, 426, 437, 454, 467, 480, 494, 525, 541, 557, 577]. Maximum [100, 401, 469, 495, 461, 505]. May [347]. Mean [536, 170, 291]. Means [122, 17, 416, 373, 399, 428, 407]. Measurement [145, 486]. Measuring [199, 391, 383, 146]. Mechanism [307]. Median [151]. Median-Effective [151]. Medical [141, 37]. melanogaster [472]. Members [329]. Memoriam [93]. Metameter [6, 1]. Method [257, 321, 457, 545, 549, 195, 542, 387, 499, 375, 526, 529, 132, 414, 159, 461, 359]. Methode [499]. Methodology [482]. Methods [33, 31, 100, 167, 32, 357, 87, 444, 455, 120, 516, 264]. Micro [198]. Micro-Organisms [198]. Misclassification [267]. Missing [447, 122, 433, 553, 536, 259, 286, 409, 367, 342, 475, 246]. Missing-Plot [259]. Mitscherlich [214]. Mixed [378, 338, 342, 573]. Mixed-Up [342]. Mixtures [208, 566]. Model [501, 571, 257, 310, 344, 438, 512, 547, 256]. Moments [233]. Mood [504]. Most [17]. Moving [549, 123]. Moving-Average [123]. Multi [356, 184, 157]. Multi-Hit [356, 184]. Multi-Stimulus [157]. Multinomial [440]. Multinomially [496]. Multiple [295, 487, 156, 341, 400, 279, 416, 553, 443, 373, 399, 240, 563, 578]. Multiple-Decision [487]. Multivariate [345, 48, 462, 37].

Natural [512, 468, 178]. Nature [427]. Near [520]. Near-Rotatable [520]. Negative [182, 183]. Net [87]. News [83, 116, 223, 329, 379, 394, 411, 425, 436, 452, 466, 479, 492, 508, 524, 556, 575]. Neyman [202, 293]. No [518]. Non [560, 520, 566, 513, 521]. Non-Center [520]. Non-Independent [521]. Non-Interactive [566]. Non-Replicated [513]. Non-Significant [560]. Normal [487, 213]. Note [266, 328, 341, 372, 183, 356, 387, 260, 173, 259, 558, 286, 318, 39, 369, 488, 503, 536, 535, 568, 444, 445, 489, 409, 463]. Notes [83, 116, 87]. nouvelles [216]. NPH [208]. Nuage [311]. nubilalis [404, 489]. Number [158, 17, 85, 362, 413, 548]. Numbers [320, 373, 74]. Numerical [357].

Obituary [493]. Observations [448, 229, 66, 550, 271, 74]. Observed [91]. Obtaining [486]. Occurring [197]. Odor [3]. Offspring [172]. One [487, 488, 347]. Operations [233]. Optimum [339, 558, 543, 439]. Organic [169]. Organisms [198, 242]. Orthogonal [534, 476, 526]. Other [54]. Outliers [171]. Outline [53].

P [93]. Pain [482]. Pair [544, 281]. Paired [268, 383, 257, 403, 280]. Pairs [371]. Paper [11]. Papers [57, 334, 446, 447, 448, 505]. Parabola [402]. Parallel [414]. paramecium [256]. Parameter [293]. Parameters [30, 292, 444, 528]. Parent [172]. Partially [101, 419, 86, 135, 316, 246]. Partition [309, 440, 265]. Past [542]. Path [360]. Paths [530]. Pearson [213]. Peck [53]. Percentage [244]. Perennial [376, 296]. Period [519]. Peritoneal [241]. Pharmacology [145]. Phase [514, 308]. Physiological [390]. pieces [573]. Plaid [181]. Plan [355]. Planning [65]. Plans [141]. Plant [28]. Plantations [269]. Plate [40]. Pleiotropism [559]. Plot [447, 433, 553, 471, 181, 259, 409, 367, 340]. Point [535]. Points [520, 311]. Poisson [155, 314, 4, 536, 260, 212, 370, 461]. Poliomyelitis [333]. Polymorphism [418]. Polynomials [526]. Population [185, 362, 360, 420, 292, 36, 468, 178]. Populations [487, 384, 359]. Posteriori [499]. Potency [341, 56, 240, 414]. Pots [422]. Power [7, 548]. Precision [391, 459, 415, 78]. Prediction [284]. Prefer [112]. Preference [565, 310]. Preparation [51]. Preparations [158]. Preparing [76]. Presence [498, 545]. Present [10, 71, 542]. Presidential [333, 216]. Prime [7]. Prime-Power [7]. Priori [499]. Probability [310, 470]. Probable [17]. Probit [65]. Probits [112]. Problem [143, 142, 133]. problèmes [102]. Problems [199, 60, 168, 146, 376, 102]. Procedure [487, 3, 40]. Proceedings [215]. Processes [446, 405, 19, 194]. Processing [171]. Product [400]. Production [239]. Products [136]. Progeny [558, 439, 548, 407]. Program [217, 331]. Programming [355]. Proneness [113]. Prophylactic [283]. Prophylaxis [283]. Proportionate [74]. Proportions [313, 371, 119, 516]. Proteins [54]. Psychiatry [546]. Pulmonary [501]. Purpose [355]. Purposes [132]. Pyrausta [404, 489].

Quadratic [567]. Quadruple [515]. Quail [569]. Quality [167, 357, 203]. Quantal [564, 211, 323, 502, 346, 566, 415, 496]. Quantal-Response [502]. Quantifying [3]. Quantitative [283, 284]. Quantity [203]. quasi [368]. quasi-Latin [368]. Query [24, 220, 22, 25, 41, 42, 43, 44, 45, 80, 504, 67, 520, 234, 302, 14, 553, 114, 300, 476, 235, 325, 569, 79, 219, 422, 128, 349, 273, 152, 348, 261, 94, 95, 106, 127, 137, 160, 161, 221, 174, 175, 187, 188, 189, 206, 249, 250, 377, 392, 408, 2, 15, 262, 287, 537, 521, 23, 301, 363].

Rahn [85]. Random [210, 344, 532]. Randomised [500]. Randomize [110]. Randomized [122, 270, 285, 413]. Range [279, 416, 443, 373, 399]. Rank [257, 372, 258, 563]. Rankit [383]. Ranks [133, 121]. Rapid [568]. Rat [241]. Rate [570, 517]. Rates [443, 87, 516, 358]. Ratio [295, 534, 158]. Ratios [354]. Re [85]. Re-Examination [85]. Reaction [417]. Recherche [216]. Reconsideration [482]. Records [527]. Recovery [398]. Rectangular [104, 515, 5, 126, 205, 86, 135, 173, 101]. Rectiligne [311]. Rectilinear [311]. Red [441]. Reference [50, 292]. Registration [218, 332]. Regression [400, 214, 299, 172, 420, 375, 99, 73, 530, 537]. Regressions [535, 428]. Related [294]. Relation [72, 346, 239, 62]. Relationship [471]. Relationships [91]. Relative [341, 472, 240, 61, 317, 358]. Relatives [255]. Relief [482]. Remarks [36]. Removal [197, 359]. Repeatability [271]. Repeated [420]. Repellant [195]. Replicate [531]. Replicated [513, 414, 316]. Replication [520, 196, 378, 338]. Replications [553, 373]. Reproduction [87]. Research [216, 35, 132]. Respect [66]. Response [6, 199, 145, 227, 307, 1, 502, 567, 239, 415, 146, 198, 240, 120, 157, 272, 496]. Response-Time [120, 157, 272]. Responses [564, 92, 346, 566]. Restricted [511]. Results [258, 551]. Reticulocyte [61]. retortaeformis [210]. Ridit [456]. Right [53]. Roadside [569]. Roger [493]. Rotatable [520]. Rotation [513]. Rotations [254]. Routine [100, 445]. Row [233, 398]. Rules [291].

Sample [448, 229, 156, 443, 77, 536, 532, 370, 578]. Sampling [76, 209, 285, 87, 8, 413, 552, 40, 460, 474, 102]. Sanitary [18]. Scales [21]. Schedules [76]. Scientific [217, 331]. Scope [372]. Scores [30, 322, 417]. Scoring [100, 166]. Sea [470, 460]. Selecting [487]. Selection [29, 59, 511, 512, 543, 439, 256]. Self [384]. Semi [564]. Semi-Quantal [564]. Seminar [352]. Sensitivities [571, 442, 547]. Sensitivity [271, 281]. Sensory [533]. Sequential [448, 229, 487, 371, 141, 156, 140, 8, 578]. Sera [56, 111]. Series [4, 378, 338, 369]. Setting [529]. Several [487, 516]. Sewage [19]. Sexes [185]. Shape [201]. Short [265, 309]. Short-Cut [265, 309]. Sib [543]. Sibs [548]. Significance [371, 4, 568, 239, 370, 340, 121, 521]. Significant [560]. Significantly [560]. Similar [571, 442, 547]. Simple [38, 550, 444, 135, 375, 526]. Simplex [294]. Simplified [374, 382, 512]. Simultaneous [343, 111]. Single [513]. Singly [200, 213, 180, 382]. Sire [548]. Size [448, 229, 38, 471, 558, 543, 439, 317]. Sizes [104, 443]. Slope [295, 534, 158]. Slope-Ratio [158]. Small [370]. Smallest [567]. Society [481, 27, 46, 68, 82, 108, 130, 163, 176, 192, 222, 237, 275, 289, 335, 351, 365, 395, 410, 424, 435, 451, 465, 478, 491, 507, 523, 539, 555, 574]. Soil [318, 248]. Solubility [100]. Solution [344, 499]. Solutions [463]. Some [66, 418, 227, 167, 264, 105, 500, 269, 258, 271, 403, 444, 168, 36, 146, 376, 370, 455, 516, 125, 551, 246, 199, 550]. Sorting [533]. Spaced [526]. Sparse [61]. Special [50]. Specific [358]. Specifying [354]. Sperm [472]. Split [181, 340]. Split-Plot [181, 340]. Sprays [383, 195]. Square [560, 518, 135, 551, 368]. Squares [504, 136, 181, 88, 291, 39, 316]. Standard [484, 51]. Standards [50]. States [333]. Statistic [386]. Statistical [199, 295, 197, 167, 546, 158, 18, 19, 381, 403, 87, 168, 146, 455, 111, 97]. Statistics [113, 216, 90, 474]. Statistique [216, 90]. Status [10, 71]. Sterilised [468]. Stimuli [502]. Stimulus [157]. Stochastic [501, 446, 405, 562, 255, 194, 256]. Straight [387]. Stratification [361, 432, 319]. Strawberry [204]. Strength [438]. strengthening [264]. Students [297]. Studies [565, 446, 405, 18, 283, 482, 239, 21]. Study [294, 550, 384, 390, 62]. Sub [320, 74]. Sub-Class [320]. Sub-Classes [74]. Subject [527, 387]. Subjective [199, 145, 3, 357, 146]. Substances [54]. Successive [269]. Suggested [233]. Sulfamerazine [89]. Sum [563]. Sums [136]. Surface [307]. Surfaces [227, 307, 567]. Survey [47]. Surveys [389]. Survival [356, 470, 184, 272]. Survivor [322]. Susceptibility [458]. Symmetrical [158]. Symmetry [37]. Symptoms [322]. System [307]. Systematic [532, 88]. Systems [546].

Table [504, 321, 103]. Tables [565, 309, 401, 213, 372, 243, 265, 568, 123, 342, 151, 267]. Taste [167, 502, 258, 271, 168]. Taste-Stimuli [502]. Tasting [310]. Teaching [13, 12]. Technique [433, 569, 376, 460]. Techniques [447, 19, 403, 367, 28]. Temps [55]. Terry [310]. Test [266, 158, 38, 553, 258, 518, 568, 563, 121]. Testing [100, 167, 72, 457, 545, 439, 529]. Tests [313, 279, 416, 63, 443, 271, 281, 373, 399, 98, 168, 558, 370, 91, 408, 340, 37, 551, 264]. Tetraploids [512]. Textile [78]. Texture [3]. Their [92, 239, 151]. Theoretical [384]. Theory [54, 7, 502, 566, 280, 53, 298]. there [122]. Third [191, 51]. Three [518]. Three-Factor [518]. Thumb [291]. Thyroid [111]. Thyrotrophic [111]. Ties [544]. Time [488, 120, 157, 272, 55]. Times [322, 496]. Titrations [415]. Tomato [21]. Tool [77]. Toxoplasma [56]. Transects [485]. Transformation [211]. Transformations [231, 232, 21]. Treated [532]. Treatment [228, 19, 553, 545, 428, 97]. Treatments [498, 500, 563, 121]. Tree [285]. Trend [38]. Trends [313, 197, 527]. Tri [344]. Tri-Allelic [344]. Triad [271, 281]. Trial [333]. Trials [204, 371]. Triangular [200, 321, 180]. Trichostrongylus [210]. Triple [205]. Trout [89]. Truncated [213, 155, 314, 260, 212]. Truncation [272]. Tsetse [468]. Tuberculosis [501]. Two [122, 498, 514, 293, 253, 166, 417, 535, 320, 245, 308, 531, 74, 121, 521]. Two-Way [320, 245, 74]. Type [293, 414]. Types [47, 472].

Unbalanced [430]. Unbiased [561]. Underlying [485]. Unequal [104, 476, 320, 373, 74]. Unequally [526]. Unified [566]. Uniformity [204].

Unintentional [381]. Uniqueness [169]. Unit [498]. United [333]. units [256]. Unknown [487]. Urea [91]. Use [565, 304, 487, 456, 76, 208, 294, 247, 214, 562, 196, 255, 88, 21, 485, 406, 548, 151, 121]. Used [403]. Uses [427, 35]. Using [104, 322, 415, 549].

Vaccination [239]. Vaccine [333]. Values [122, 259, 475]. Varenna [352]. Variability [510, 471]. Variable [526]. Variables [417, 387]. Variance [571, 487, 20, 76, 209, 75, 71, 430, 231, 232, 270, 354, 355, 243, 186, 77, 459, 320, 420, 292, 444, 445, 170, 559, 312, 552, 386, 291, 547, 74, 484, 407]. Variances [122, 72]. Variation [282, 497, 483, 385, 178]. Variations [124]. Various [487, 231, 232]. Varying [488]. Vectors [440]. Veneer [357]. versus [563]. Viabilities [512]. Visit [457]. Volume [69, 117, 164, 224, 276, 336, 396, 453, 509, 576].

W. [11]. Wald [387]. Walk [210]. Water [470]. Way [320, 245, 74]. Wear [60]. Weather [299]. Weight [62, 21]. Weighted [170]. Weighting [209]. Which [148, 228]. White [124]. White-Cell [124]. Whitefish [8]. Wilcoxon [134]. Winsor [93]. without [544]. Wool [76].

Yield [204]. Yields [471].

Zero [536].

References

Finney:1949:CRM

[1] D. J. Finney. The choice of a response metameter in bio-assay. Biometrics, 5(4):261–272, December 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 3001511. See correction [6].

Tukey:1949:Q

[2] John W. Tukey. 74. Query. Biometrics, 5(4):337–340, December 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

     https://www.jstor.org/stable/3001517. See correction [24].

Hopkins:1950:PQS

[3] J. W. Hopkins. A procedure for quantifying subjective appraisals of odor, flavor and texture of foodstuffs. Biometrics, 6(1):1–16, March 1950. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

     https://www.jstor.org/stable/3001419.

Fisher:1950:SDE

[4] R. A. Fisher. The significance of deviations from expectation in a Poisson series. Biometrics, 6(1):17–24, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001420.

Grundy:1950:EER

[5] P. M. Grundy. The estimation of error in rectangular lattices. Biometrics, 6(1):25–33, March 1950. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001421.

Anonymous:1950:CCR

[6] Anonymous. Correction: The choice of a response metameter in bio-assay. Biometrics, 6(1):33, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 3001422. See [1].

Federer:1950:GTP

[7] W. T. Federer. The general theory of prime-power lattice designs. Biometrics, 6(1):34–58, March 1950. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001423.

Oakland:1950:ASA

[8] G. B. Oakland. An application of sequential analysis to whitefish sampling. Biometrics, 6(1):59–67, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001424.

Anonymous:1950:I

[9] Anonymous. [Introduction]. Biometrics, 6(1):69–74, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001425.

Cochran:1950:PSB

[10] William G. Cochran. The present status of biometry. Biometrics, 6(1): 75–78, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001426. See

discussion [11].

Anonymous:1950:DFP

[11] Anonymous. Discussion following paper by W. G. Cochran. Biometrics, 6 (1):79–84, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001427. See

[10].

Bartlett:1950:TEB

[12] M. S. Bartlett. Teaching and education in biometry. Biometrics, 6(1):85– 95, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001428.

Anonymous:1950:DFT

[13] Anonymous. Discussion following teaching and education in biometry. Biometrics, 6(1):96–98, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001429.

Finney:1950:Q

[14] D. J. Finney. 75. Query. Biometrics, 6(1):99–102, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001430.

Tukey:1950:Q

[15] John W. Tukey. 76. Query. Biometrics, 6(1):102, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001431.

Anonymous:1950:FMa

[16] Anonymous. Front matter. Biometrics, 6(1):??, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001418.

Cochran:1950:EBD

[17] William G. Cochran. Estimation of bacterial densities by means of the “Most probable number”. Biometrics, 6(2):105–116, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001491.

Eliassen:1950:SAS

[18] Rolf Eliassen. Statistical analysis in sanitary engineering laboratory studies. Biometrics, 6(2):117–126, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001492.

Fertig:1950:AST

[19] John W. Fertig and Austin N. Heller. The application of statistical techniques to sewage treatment processes. Biometrics, 6(2):127–135, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001493.

Bross:1950:FIV

[20] Irwin Bross. Fiducial intervals for variance components. Biometrics, 6(2): 136–144, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001494.

Powers:1950:DSU

[21] LeRoy Powers. Determining scales and the use of transformations in studies on weight per locule of tomato fruit. Biometrics, 6(2):145–163, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001495.

Anonymous:1950:Qa

[22] Anonymous. 77. Query. Biometrics, 6(2):164–166, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001496.

Winsor:1950:Q

[23] C. P. Winsor. 78. Query. Biometrics, 6(2):167–168, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001497.

Anonymous:1950:CQ

[24] Anonymous. Correction: In query 74. Biometrics, 6(2):168, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001499. See [2].

Anonymous:1950:Qb

[25] Anonymous. 79. Query. Biometrics, 6(2):168, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001498.

Anonymous:1950:A

[26] Anonymous. Abstracts. Biometrics, 6(2):169–194, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001500.

Anonymous:1950:BSa

[27] Anonymous. The Biometric Society. Biometrics, 6(2):195–199, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001501.

Yates:1950:ETP

[28] Frank Yates. Experimental techniques in plant improvement. Biometrics, 6(2):200–207, June 1950. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001502.

Cavalli:1950:ASC

[29] Luigi L. Cavalli. The analysis of selection curves. Biometrics, 6(2):208– 220, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001503.

Finney:1950:SEP

[30] D. J. Finney. Scores for the estimation of parameters. Biometrics, 6(2): 221–225, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001504.

Anonymous:1950:DRA

[31] Anonymous. Discussion on recent applications of biometrical methods in genetics. Biometrics, 6(2):226–227, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001505.

Davies:1950:BMC

[32] O. L. Davies. Biometric methods in the chemical industry. Biometrics, 6 (2):228, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001506.

Anonymous:1950:DFB

[33] Anonymous. Discussion following biometric methods in the chemical industry. Biometrics, 6(2):229–230, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001507.

Anonymous:1950:FMb

[34] Anonymous. Front matter. Biometrics, 6(2):??, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001490.

Davies:1950:CUF

[35] O. L. Davies and W. A. Hay. The construction and uses of fractional factorial designs in industrial research. Biometrics, 6(3):233–249, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001821.

Moran:1950:SRA

[36] P. A. P. Moran. Some remarks on animal population dynamics. Biometrics, 6(3):250–258, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001822.

Votaw:1950:CST

[37] D. F. Votaw, Jr., A. W. Kimball, and J. A. Rafferty. Compound symmetry tests in the multivariate analysis of medical experiments. Biometrics, 6(3): 259–281, September 1950. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001823.

Elfving:1950:STT

[38] G. Elfving and J. H. Whitlock. A simple trend test with application to erythrocyte size data. Biometrics, 6(3):282–288, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001824.

Youden:1950:NFF

[39] W. J. Youden. A note on the four by four Latin squares. Biometrics, 6(3): 289–290, September 1950. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001825.

Sandelius:1950:ISP

[40] Martin Sandelius. An inverse sampling procedure for bacterial plate counts. Biometrics, 6(3):291–292, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001826.

Anonymous:1950:Qc

[41] Anonymous. 80. Query. Biometrics, 6(3):293–294, September 1950. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001827.

Anonymous:1950:Qd

[42] Anonymous. 81. Query. Biometrics, 6(3):294–295, September 1950. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001828.

Anonymous:1950:Qe

[43] Anonymous. 82. Query. Biometrics, 6(3):295–296, September 1950. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001829.

Anonymous:1950:Qf

[44] Anonymous. 83. Query. Biometrics, 6(3):296–297, September 1950. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001830.

Anonymous:1950:Qg

[45] Anonymous. 84. Query. Biometrics, 6(3):298–299, September 1950. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001831.

Anonymous:1950:BSb

[46] Anonymous. The Biometric Society. Biometrics, 6(3):300, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001832.

Cox:1950:STE

[47] Gertrude M. Cox. A survey of types of experimental designs. Biometrics, 6 (3):301–302, September 1950. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001833.

Quenouill:1950:ME

[48] M. H. Quenouill. Multivariate experimentation. Biometrics, 6(3):303–316, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001834.

Anonymous:1950:DED

[49] Anonymous. Discussion on experimental design. Biometrics, 6(3):317– 319, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001835.

Irwin:1950:BAS

[50] J. O. Irwin. Biological assays with special reference to biological standards. Biometrics, 6(3):320–321, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001836.

Perry:1950:ACA

[51] W. L. M. Perry. The analysis of a collaborative assay of the Third International Digitalis Standard preparation. Biometrics, 6(3):322–324, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/3001837.

Anonymous:1950:DBA

[52] Anonymous. Discussion on biometrical aspects of biological asay. Biometrics, 6(3):325–329, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001838.

Rapoport:1950:OMT

[53] Anatol Rapoport. Outline of a mathematical theory of peck right. Biometrics, 6(3):330–341, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001839.

Boeri:1950:MTE

[54] Enzo Boeri. On the mathematical theory of the equilibrium of interactions between proteins and other substances. Biometrics, 6(3):342–343, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001840.

Schwartz:1950:CLF

[55] Daniel Schwartz, Jean Cuzin, and André Renier. Contribution à l’Analyse factorielle d’un temps d’Incubation. (French) [Contribution to the factorial analysis of an incubation time]. Biometrics, 6(3):344, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001841.

Nass:1950:FHC

[56] Ch. A. G. Nass. Free hand curves in estimating the potency of human sera against toxoplasma. Biometrics, 6(3):345–346, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001842.

Anonymous:1950:DCP

[57] Anonymous. Discussion on contributed papers. Biometrics, 6(3):347, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001843.

Anonymous:1950:FMc

[58] Anonymous. Front matter. Biometrics, 6(3):??, September 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001820.

Fisher:1950:GFC

[59] R. A. Fisher. Gene frequencies in a cline determined by selection and diffusion. Biometrics, 6(4):353–361, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001780.

Box:1950:PAG

[60] G. E. P. Box. Problems in the analysis of growth and wear curves. Biometrics, 6(4):362–389, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001781.

Schneiderman:1950:RFS

[61] Marvin Schneiderman and George Brecher. The relative frequency of sparse cell elements — an application to reticulocyte blood counts. Biometrics, 6(4):390–394, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/3001782.

Norval:1950:BSE

[62] Mildred A. Norval and Nancy King. A biometric study of the excretion of corticosteroids in children in relation to age, height and weight. Biometrics, 6(4):395–398, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001783.

Greenhouse:1950:EDT

[63] Samuel W. Greenhouse and Nathan Mantel. The evaluation of diagnostic tests. Biometrics, 6(4):399–412, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001784.

Bross:1950:ELC

[64] Irwin Bross. Estimates of the LD 50: a critique. Biometrics, 6(4):413– 423, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001785. See

comments [105].

Healy:1950:PPA

[65] M. J. R. Healy. The planning of probit assays. Biometrics, 6(4):424–431, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001786.

Berkson:1950:SOR

[66] Joseph Berkson. Some observations with respect to the error of bio-assay. Biometrics, 6(4):432–434, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001787.

Bliss:1950:Q

[67] C. I. Bliss and C. P. Winsor. 85. Query. Biometrics, 6(4):435–437, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001788.

Anonymous:1950:BSc

[68] Anonymous. The Biometric Society. Biometrics, 6(4):438–440, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001789.

Anonymous:1950:VI

[69] Anonymous. Volume information. Biometrics, 6(4):441–447, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001778.

Anonymous:1950:FMd

[70] Anonymous. Front matter. Biometrics, 6(4):??, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001779.

Crump:1951:PSV

[71] S. Lee Crump. The present status of variance component analysis. Biometrics, 7(1):1–16, March 1951. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001600.

Cochran:1951:TLR

[72] W. G. Cochran. Testing a linear relation among variances. Biometrics, 7(1):17–32, March 1951. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001601.

Tukey:1951:CR

[73] John W. Tukey. Components in regression. Biometrics, 7(1):33–69, March 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001602.

Smith:1951:AVU

[74] H. Fairfield Smith. Analysis of variance with unequal but proportionate numbers of observations in the sub-classes of a two-way classification. Biometrics, 7(1):70–74, March 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001603.

Comstock:1951:CEV

[75] R. E. Comstock and H. F. Robinson. Consistency of estimates of variance components. Biometrics, 7(1):75–82, March 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001604.

Cameron:1951:UCV

[76] J. M. Cameron. The use of components of variance in preparing schedules for sampling of baled wool. Biometrics, 7(1):83–96, March 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001605.

Hendricks:1951:VCT

[77] Walter A. Hendricks. Variance components as a tool for the analysis of sample data. Biometrics, 7(1):97–101, March 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001606.

Whitwell:1951:EPT

[78] John C. Whitwell. Estimating precision of textile instruments. Biometrics, 7(1):102–112, March 1951. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001607.

Johnson:1951:Q

[79] Palmer O. Johnson. 86. Query. Biometrics, 7(1):113–114, March 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001608.

Anonymous:1951:Q

[80] Anonymous. 87. Query. Biometrics, 7(1):114–116, March 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001609.

Anonymous:1951:Aa

[81] Anonymous. Abstracts. Biometrics, 7(1):117–126, March 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001610.

Anonymous:1951:BSa

[82] Anonymous. The Biometric Society. Biometrics, 7(1):127–129, March 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001611.

Anonymous:1951:NNa

[83] Anonymous. News and notes. Biometrics, 7(1):130–131, March 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001612.

Anonymous:1951:FMa

[84] Anonymous. Front matter. Biometrics, 7(1):??, March 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001599.

Finney:1951:RER

[85] D. J. Finney and L. Martin. A re-examination of Rahn’s data on the number of genes in bacteria. Biometrics, 7(2):133–144, June 1951. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001741.

Nair:1951:RLP

[86] K. R. Nair. Rectangular lattices and partially balanced incomplete block designs. Biometrics, 7(2):145–154, June 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/3001742. See corrigenda [101].

Koop:1951:NEG

[87] J. C. Koop. Notes on the estimation of gross and net reproduction rates by methods of statistical sampling. Biometrics, 7(2):155–166, June 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001743.

Nissen:1951:USS

[88] Øivind Nissen. The use of systematic 5 × 5 squares. Biometrics, 7(2):167– 170, June 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001744.

Gutsell:1951:ESE

[89] James S. Gutsell. The effect of sulfamerazine on the erythrocyte and hemoglobin content of trout blood. Biometrics, 7(2):171–179, June 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001745.

Frechet:1951:ASM

[90] Maurice Fréchet. Sur une application de la statistique mathématique à la biologie. (French) [On an application of mathematical statistics to biology]. Biometrics, 7(2):180–184, June 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001746.

Smith:1951:RAC

[91] H. Fairfield Smith. Relationships among characters observed in urea clearance tests. Biometrics, 7(2):185–199, June 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001747.

Gridgeman:1951:EBA

[92] Norman T. Gridgeman. On the errors of biological assays with graded responses, and their graphical derivation. Biometrics, 7(2):200–221, June 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001748.

Anonymous:1951:MCP

[93] Anonymous. In memoriam: Charles P. Winsor. Biometrics, 7(2):221, June 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001749.

Snedecor:1951:Qa

[94] George W. Snedecor. 88. Query. Biometrics, 7(2):222, June 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001750.

Snedecor:1951:Qb

[95] George W. Snedecor. 89. Query. Biometrics, 7(2):222–223, June 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001751.

Anonymous:1951:FMb

[96] Anonymous. Front matter. Biometrics, 7(2):??, June 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001740.

vanderReyden:1951:STD

[97] D. van der Reyden. Statistical treatment of the data of blood concentration of a drug. Biometrics, 7(3):227–239, September 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001807.

Mantel:1951:ECD

[98] Nathan Mantel. Evaluation of a class of diagnostic tests. Biometrics, 7(3): 240–246, September 1951. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001808.

Stevens:1951:AR

[99] W. L. Stevens. Asymptotic regression. Biometrics, 7(3):247–267, September 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/3001809.

Bailey:1951:TSM

[100] Norman T. J. Bailey. Testing the solubility of maximum likelihood equations in the routine application of scoring methods. Biometrics, 7(3): 268–274, September 1951. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001810.

Anonymous:1951:CRL

[101] Anonymous. Corrigenda: Rectangular lattices and partially balanced incomplete block designs. Biometrics, 7(3):274, September 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001811. See [86].

Lamotte:1951:CPE

[102] M. Lamotte and M. Schutzenberger. Sur certains problèmes d’estimation dans les cas de double échantillonnage. (French) [On certain estimation problems in the case of double sampling]. Biometrics, 7(3):275–282, September 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001812.

Leslie:1951:CCT

[103] P. H. Leslie. The calculation of χ^{2} for an r×c contingency table. Biometrics, 7(3):283–286, September 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001813.

Clarke:1951:ARL

[104] Pamela M. Clarke. An analysis of rectangular lattices with unequal block sizes, using inter-block information. Biometrics, 7(3):287–294, September 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001814.

Cornfield:1951:SCE

[105] Jerome Cornfield and Nathan Mantel. Some comments on “Estimates of the LD 50: a critique”. Biometrics, 7(3):295–298, September 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001815. See [64].

Snedecor:1951:Qc

[106] George W. Snedecor. 90. Query. Biometrics, 7(3):299, September 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001816.

Anonymous:1951:Ab

[107] Anonymous. Abstracts. Biometrics, 7(3):300–302, September 1951. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001817.

Anonymous:1951:BSb

[108] Anonymous. The Biometric Society. Biometrics, 7(3):303–305, September 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001818.

Anonymous:1951:FMc

[109] Anonymous. Front matter. Biometrics, 7(3):??, September 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001806.

Greenberg:1951:WR

[110] B. G. Greenberg. Why randomize? Biometrics, 7(4):309–322, December 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001653.

Steinmetz:1951:SAS

[111] C. H. Steinmetz. Statistical aspects of the simultaneous detection of thyroid and thyrotrophic hormones in human sera, based on the data of D’Angelo and Gordon. Biometrics, 7(4):323–326, December 1951. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001654.

Berkson:1951:WPL

[112] Joseph Berkson. Why I prefer logits to probits. Biometrics, 7(4):327–339, December 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001655.

Arbous:1951:ASC

[113] A. G. Arbous and J. E. Kerrich. Accident statistics and the concept of accident-proneness. Biometrics, 7(4):340–432, December 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001656.

Fisher:1951:Q

[114] R. A. Fisher. 91. Query. Biometrics, 7(4):433–434, December 1951. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001657.

Anonymous:1951:Ac

[115] Anonymous. Abstracts. Biometrics, 7(4):435, December 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001658.

Anonymous:1951:NNb

[116] Anonymous. News and notes. Biometrics, 7(4):436–452, December 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001659.

Anonymous:1951:VI

[117] Anonymous. Volume information. Biometrics, 7(4):453–459, December 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001651.

Anonymous:1951:FMd

[118] Anonymous. Front matter. Biometrics, 7(4):??, December 1951. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001652.

Dyke:1952:AFA

[119] G. V. Dyke and H. D. Patterson. Analysis of factorial arrangements when the data are proportions. Biometrics, 8(1):1–12, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001521.

Sampford:1952:ERTa

[120] M. R. Sampford. The estimation of response-time distributions. I. Fundamental concepts and general methods. Biometrics, 8(1):13–32, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001522.

White:1952:URT

[121] Colin White. The use of ranks in a test of significance for comparing two treatments. Biometrics, 8(1):33–41, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001523.

Baten:1952:VDB

[122] William Dowell Baten. Variances of differences between means when there are two missing values in randomized block designs. Biometrics, 8(1):42– 50, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001524.

Thompson:1952:CTM

[123] William R. Thompson and Carrol S. Weil. On the construction of tables for moving-average interpolation. Biometrics, 8(1):51–54, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001525.

Chamberlain:1952:EVW

[124] A. C. Chamberlain and F. M. Turner. Errors and variations in white-cell counts. Biometrics, 8(1):55–65, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001526.

Shrikhande:1952:DSB

[125] S. S. Shrikhande. On the dual of some balanced incomplete block designs. Biometrics, 8(1):66–72, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001527.

Harshbarger:1952:LRL

[126] Boyd Harshbarger and Lyle L. Davis. Latinized rectangular lattices. Biometrics, 8(1):73–84, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001528.

Snedecor:1952:Qa

[127] George W. Snedecor. 92. Query. Biometrics, 8(1):85–86, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001529.

Kimball:1952:Q

[128] A. W. Kimball. 93. Query. Biometrics, 8(1):86–88, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001530.

Anonymous:1952:Aa

[129] Anonymous. Abstracts. Biometrics, 8(1):89–96, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001531.

Anonymous:1952:BSa

[130] Anonymous. The Biometric Society. Biometrics, 8(1):97–99, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001532.

Anonymous:1952:FMa

[131] Anonymous. Front matter. Biometrics, 8(1):??, March 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001520.

Sharp:1952:MEI

[132] J. Sharp, E. L. Feinmann, and John F. Wilkinson. A method of enumeration of individual marrow elements for research purposes. Biometrics, 8(2):105–111, June 1952. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001926.

Keeping:1952:PBR

[133] E. S. Keeping. The problem of birth ranks. Biometrics, 8(2):112–119, June 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001927.

Eisenberg:1952:CLW

[134] Frieda Faiman Eisenberg. A comparison of Litchfield–Wilcoxon and Bliss estimates. Biometrics, 8(2):120–121, June 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001928.

Nair:1952:APB

[135] K. R. Nair. Analysis of partially balanced incomplete block designs illustrated on the simple square and rectangular lattices. Biometrics, 8(2): 122–155, June 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001929.

Hammersley:1952:CSS

[136] J. M. Hammersley. The computation of sums of squares and products on a desk calculator. Biometrics, 8(2):156–168, June 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001930.

Snedecor:1952:Qb

[137] George W. Snedecor. 94. Query. Biometrics, 8(2):169–171, June 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001931.

Anonymous:1952:Ab

[138] Anonymous. Abstracts. Biometrics, 8(2):172–179, June 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001932.

Anonymous:1952:FMb

[139] Anonymous. Front matter. Biometrics, 8(2):??, June 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001925.

Fisher:1952:SE

[140] R. A. Fisher. Sequential experimentation. Biometrics, 8(3):183–187, September 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001546.

Bross:1952:SMP

[141] I. Bross. Sequential medical plans. Biometrics, 8(3):188–205, September 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001547.

Izzo:1952:CP

[142] J. L. Izzo. The clinical problem. Biometrics, 8(3):206–209, September 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001548.

Crump:1952:BAP

[143] S. Lee Crump. The biometrical aspect of the problem. Biometrics, 8(3): 210–215, September 1952. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001549.

Anonymous:1952:Da

[144] Anonymous. Discussion. Biometrics, 8(3):215–217, September 1952. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001550.

Beecher:1952:EPM

[145] Henry K. Beecher. Experimental pharmacology and measurement of the subjective response. Biometrics, 8(3):218–220, September 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001551.

Mosteller:1952:SSP

[146] Frederick Mosteller. Some statistical problems in measuring the subjective response to drugs. Biometrics, 8(3):220–226, September 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001552. See correction [199].

Anonymous:1952:Db

[147] Anonymous. Discussion. Biometrics, 8(3):227–231, September 1952. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001553.

Greiner:1952:BCW

[148] Theodore Greiner and Harry Gold. Biologic considerations which determine the design. Biometrics, 8(3):232–237, September 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001554.

Bliss:1952:EEC

[149] C. I. Bliss. Estimation of the error in clinical assay. Biometrics, 8(3): 237–245, September 1952. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001555.

Anonymous:1952:Dc

[150] Anonymous. Discussion. Biometrics, 8(3):245–248, September 1952. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001556.

Weil:1952:TCC

[151] Carrol S. Weil. Tables for convenient calculation of median-effective dose (LD 50 or ED 50) and instructions in their use. Biometrics, 8(3):249–263, September 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001557.

Ostle:1952:Q

[152] Bernard Ostle. 95. Query. Biometrics, 8(3):264–266, September 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001558.

Anonymous:1952:Ac

[153] Anonymous. Abstracts. Biometrics, 8(3):267–272, September 1952. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001559.

Anonymous:1952:FMc

[154] Anonymous. Front matter. Biometrics, 8(3):??, September 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001545.

David:1952:TP

[155] F. N. David and N. L. Johnson. The truncated Poisson. Biometrics, 8(4): 275–285, December 1952. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001863.

Chapman:1952:IMS

[156] Douglas G. Chapman. Inverse, multiple and sequential sample censuses. Biometrics, 8(4):286–306, December 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001864. See erratum [578].

Sampford:1952:ERTb

[157] M. R. Sampford. The estimation of response-time distributions. II. Multistimulus distributions. Biometrics, 8(4):307–369, December 1952. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001865.

Clarke:1952:SAS

[158] Pamela M. Clarke. Statistical analysis of symmetrical slope-ratio assays of any number of test preparations. Biometrics, 8(4):370–379, December 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001866.

Skory:1952:AMM

[159] John Skory. Automatic machine method of calculating contingency X^{2}. Biometrics, 8(4):380–382, December 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001867.

Snedecor:1952:Qc

[160] George W. Snedecor. 96. Query. Biometrics, 8(4):383–384, December 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001868.

Snedecor:1952:Qd

[161] George W. Snedecor. 97. Query. Biometrics, 8(4):384–385, December 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001869.

Anonymous:1952:Ad

[162] Anonymous. Abstracts. Biometrics, 8(4):386–391, December 1952. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001870.

Anonymous:1952:BSb

[163] Anonymous. The Biometric Society. Biometrics, 8(4):392–394, December 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001639.

Anonymous:1952:VI

[164] Anonymous. Volume information. Biometrics, 8(4):395–401, December 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001861.

Anonymous:1952:FMd

[165] Anonymous. Front matter. Biometrics, 8(4):??, December 1952. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001862.

Hopkins:1953:LFS

[166] J. W. Hopkins. Laboratory flavor scoring: Two experiments in incomplete blocks. Biometrics, 9(1):1–21, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001629.

Bradley:1953:SSM

[167] Ralph Allan Bradley. Some statistical methods in taste testing and quality evaluation. Biometrics, 9(1):22–38, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001630.

Mason:1953:SPD

[168] David D. Mason and E. James Koch. Some problems in the design and statistical analysis of taste tests. Biometrics, 9(1):39–46, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001631.

Kruskal:1953:ULO

[169] William H. Kruskal. On the uniqueness of the line of organic correlation. Biometrics, 9(1):47–58, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001632.

Meier:1953:VWM

[170] Paul Meier. Variance of a weighted mean. Biometrics, 9(1):59–73, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001633.

Dixon:1953:PDO

[171] W. J. Dixon. Processing data for outliers. Biometrics, 9(1):74–89, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001634.

Kempthorne:1953:EHR

[172] O. Kempthorne and O. B. Tandon. The estimation of heritability by regression of offspring on parent. Biometrics, 9(1):90–100, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001635.

Nair:1953:NRL

[173] K. R. Nair. A note on rectangular lattices. Biometrics, 9(1):101–106, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/3001636.

Snedecor:1953:Qa

[174] Geroge W. Snedecor. 98. Query. Biometrics, 9(1):107–108, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001637.

Snedecor:1953:Qb

[175] Geroge W. Snedecor. 99. Query. Biometrics, 9(1):108–110, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001638.

Anonymous:1953:BSa

[176] Anonymous. The Biometric Society. Biometrics, 9(1):111–113, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001639.

Anonymous:1953:FMa

[177] Anonymous. Front matter. Biometrics, 9(1):??, March 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001628.

Sturtevant:1953:AVN

[178] Frank M. Sturtevant, Jr., Lyle D. Calvin, and Orlando Park. The analysis of variation in a natural population of lady beetles. Biometrics, 9(2):117– 126, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001846.

Youden:1953:CBD

[179] W. J. Youden and W. S. Connor. The chain block design. Biometrics, 9(2):127–140, June 1953. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001847.

Nair:1953:DAT

[180] K. R. Nair. Design and analysis of triangular singly linked blocks. Biometrics, 9(2):141–156, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 3001848. See errata [200].

Jacob:1953:SPH

[181] Walter C. Jacob. Split-plot half-plaid squares for irrigation experiments. Biometrics, 9(2):157–175, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001849.

Bliss:1953:FNB

[182] C. I. Bliss and R. A. Fisher. Fitting the negative binomial distribution to biological data. Biometrics, 9(2):176–200, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001850.

Fisher:1953:NEF

[183] R. A. Fisher. Note on the efficient fitting of the negative binomial. Biometrics, 9(2):197–200, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001850.

Kimball:1953:FMH

[184] A. W. Kimball. The fitting of multi-hit survival curves. Biometrics, 9(2): 201–211, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001851.

Goodman:1953:PGS

[185] Leo A. Goodman. Population growth of the sexes. Biometrics, 9(2):212– 225, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001852.

Henderson:1953:EVC

[186] C. R. Henderson. Estimation of variance and covariance components. Biometrics, 9(2):226–252, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001853.

Snedecor:1953:Qc

[187] George W. Snedecor. 100. Query. Biometrics, 9(2):253–255, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001854. See correction [221].

Snedecor:1953:Qd

[188] George W. Snedecor. 101. Query. Biometrics, 9(2):255–256, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001855.

Snedecor:1953:Qe

[189] George W. Snedecor. 102. Query. Biometrics, 9(2):256–258, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001856.

Anonymous:1953:A

[190] Anonymous. Abstracts. Biometrics, 9(2):259–267, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001857.

Anonymous:1953:TIB

[191] Anonymous. The third international biometric conference. Biometrics, 9(2):268–270, June 1953. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001858.

Anonymous:1953:BSb

[192] Anonymous. The Biometric Society. Biometrics, 9(2):271–272, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001859.

Anonymous:1953:FMb

[193] Anonymous. Front matter. Biometrics, 9(2):??, June 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001845.

Reid:1953:SPB

[194] A. T. Reid. On stochastic processes in biology. Biometrics, 9(3):275–289, September 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001703.

Healy:1953:MCF

[195] M. J. R. Healy. A method for comparing fly-repellant sprays. Biometrics, 9 (3):290–294, September 1953. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001704.

Kempthorne:1953:EUF

[196] O. Kempthorne and R. G. Tischer. An example of the use of fractional replication. Biometrics, 9(3):295–303, September 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001705.

Box:1953:SDE

[197] G. E. P. Box and W. A. Hay. A statistical design for the efficient removal of trends occurring in a comparative experiment with an application in biological assay. Biometrics, 9(3):304–319, September 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001706.

Peto:1953:DRE

[198] S. Peto. A dose-response equation for the invasion of micro-organisms. Biometrics, 9(3):320–335, September 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001707.

Anonymous:1953:CSS

[199] Anonymous. Correction: Some statistical problems in measuring the subjective response to drugs. Biometrics, 9(3):335, September 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001708. See [146].

Anonymous:1953:EDA

[200] Anonymous. Errata: Design and analysis of triangular singly linked blocks. Biometrics, 9(3):335, September 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001709. See [180].

Shultz:1953:AES

[201] Fred T. Shultz. Analysis of egg shape of chickens. Biometrics, 9(3): 336–353, September 1953. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001710.

Beall:1953:GNC

[202] Geoffrey Beall and Richard R. Rescia. A generalization of Neyman’s contagious distributions. Biometrics, 9(3):354–386, September 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001711.

Rayner:1953:QQI

[203] A. A. Rayner. Quality × quantity interaction. Biometrics, 9(3):387–411, September 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001712.

Baker:1953:SUY

[204] G. A. Baker and R. E. Baker. Strawberry uniformity yield trials. Biometrics, 9(3):412–421, September 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001713.

Murty:1953:ATR

[205] V. N. Murty. Analysis of a triple rectangular lattice design. Biometrics, 9 (3):422–424, September 1953. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001714.

Snedecor:1953:Qf

[206] George W. Snedecor. 103. Query. Biometrics, 9(3):425–427, September 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001715.

Anonymous:1953:FMc

[207] Anonymous. Front matter. Biometrics, 9(3):??, September 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001702.

Ciminera:1953:EUE

[208] Joseph L. Ciminera and Elwood K. Wolfe. An example of the use of extended cross-over designs in the comparison of NPH insulin mixtures. Biometrics, 9(4):431–446, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001435.

Cochran:1953:SIE

[209] William G. Cochran and Sarah Porter Carroll. A sampling investigation of the efficiency of weighting inversely as the estimated variance. Biometrics, 9(4):447–459, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001436.

Broadbent:1953:RWT

[210] S. R. Broadbent and David G. Kendall. The random walk of Trichostrongylus retortaeformis. Biometrics, 9(4):460–466, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001437.

Claringbold:1953:ATQ

[211] P. J. Claringbold, J. D. Biggers, and C. W. Emmens. The angular transformation in quantal analysis. Biometrics, 9(4):467–484, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001438.

Plackett:1953:TPD

[212] R. L. Plackett. The truncated Poisson distribution. Biometrics, 9(4): 485–488, December 1953. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001439.

Cohen:1953:TPL

[213] A. C. Cohen, Jr. and John Woodward. Tables of Pearson–Lee–Fisher functions of singly truncated normal distributions. Biometrics, 9(4):489– 497, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001440.

Gomes:1953:UMR

[214] Frederico Pimentel Gomes. The use of Mitscherlich’s regression law in the analysis of experiments with fertilizers. Biometrics, 9(4):498–516, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001441.

Anonymous:1953:P

[215] Anonymous. Proceedings. Biometrics, 9(4):517–521, December 1953. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001442.

Darmois:1953:PAD

[216] Georges Darmois. Presidential address. Dignités nouvelles de la statistique dans la recherche. (French) [new dignities from statistics in research]. Biometrics, 9(4):522–524, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001443.

Anonymous:1953:SP

[217] Anonymous. Scientific program. Biometrics, 9(4):525–526, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001444.

Anonymous:1953:R

[218] Anonymous. Registration. Biometrics, 9(4):527, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001445.

Kempthorne:1953:Q

[219] O. Kempthorne. 104. Query. Biometrics, 9(4):528–533, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001446.

Anderson:1953:Q

[220] R. L. Anderson. 105. Query. Biometrics, 9(4):533–534, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001447.

Snedecor:1953:CQ

[221] George W. Snedecor. Correction for Query 100. Biometrics, 9(4):534, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001448. See [187].

Anonymous:1953:BSc

[222] Anonymous. The Biometric Society. Biometrics, 9(4):535–537, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001449.

Anonymous:1953:N

[223] Anonymous. News. Biometrics, 9(4):538–539, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001450.

Anonymous:1953:VI

[224] Anonymous. Volume information. Biometrics, 9(4):541–548, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001433.

Anonymous:1953:FMd

[225] Anonymous. Front matter. Biometrics, 9(4):??, December 1953. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001434.

Read:1954:DCE

[226] D. R. Read. The design of chemical experiments. Biometrics, 10(1):1– 15, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001662.

Box:1954:EER

[227] G. E. P. Box. The exploration and exploitation of response surfaces: Some general considerations and examples. Biometrics, 10(1):16–60, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001663.

Calvin:1954:DBI

[228] Lyle D. Calvin. Doubly balanced incomplete block designs for experiments in which the treatment effects are correlated. Biometrics, 10(1): 61–88, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001664.

Anscombe:1954:FSS

[229] F. J. Anscombe. Fixed-sample-size analysis of sequential observations. Biometrics, 10(1):89–100, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/ stable/3001665. See corrections [448].

Cochran:1954:CED

[230] William G. Cochran. The combination of estimates from different experiments. Biometrics, 10(1):101–129, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001666.

Fisher:1954:AVV

[231] Ronald Fisher. The analysis of variance with various binomial transformations. Biometrics, 10(1):130–139; discussion, 140–151, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001667.

Fisher:1954:DAV

[232] Ronald Fisher. Discussion of the analysis of variance with various binomial transformations. Biometrics, 10(1):140–151, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001668.

Hemphill:1954:SDC

[233] F. M. Hemphill. Suggested desk calculator operations for computing moments by the row. Biometrics, 10(1):152–154, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001669.

Cochran:1954:Q

[234] W. G. Cochran. 106. Query. Biometrics, 10(1):155–157, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001670.

Gurland:1954:Q

[235] John Gurland. 107. Query. Biometrics, 10(1):157–159, March 1954. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001671.

Anonymous:1954:Aa

[236] Anonymous. Abstracts. Biometrics, 10(1):160–188, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001672.

Anonymous:1954:BSa

[237] Anonymous. The Biometric Society. Biometrics, 10(1):189–191, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001673.

Anonymous:1954:FMa

[238] Anonymous. Front matter. Biometrics, 10(1):??, March 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001661.

Meyer:1954:FSS

[239] Sven Nissen Meyer and Michael Weis Bentzon. Further studies on the significance of family factors for the response to BCG vaccination. The development of local vaccination lesions and their relation to allergy production. Biometrics, 10(2):195–207, June 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001873.

Rao:1954:ERP

[240] C. Radhakrishna Rao. Estimation of relative potency from multiple response data. Biometrics, 10(2):208–220, June 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001874.

Dews:1954:EDE

[241] Peter B. Dews, George M. Higgins, and Joseph Berkson. Error of the determination of the eosinophil count in peritoneal fluid of the rat. Biometrics, 10(2):221–226, June 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001875.

Worcester:1954:HMO

[242] Jane Worcester. How many organisms? Biometrics, 10(2):227–234, June 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001876.

Hayman:1954:AVD

[243] B. I. Hayman. The analysis of variance of diallel tables. Biometrics, 10(2):235–244, June 1954. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001877.

Bross:1954:CIP

[244] Irwin Bross. A confidence interval for a percentage increase. Biometrics, 10(2):245–250, June 1954. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001878.

Mandel:1954:CBD

[245] John Mandel. Chain block designs with two-way elimination of heterogeneity. Biometrics, 10(2):251–272, June 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001879.

Zelen:1954:ASP

[246] Marvin Zelen. Analysis for some partially balanced incomplete block designs having a missing block. Biometrics, 10(2):273–281, June 1954. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001880.

Federer:1954:UCC

[247] W. T. Federer and C. S. Schlottfeldt. The use of covariance to control gradients in experiments. Biometrics, 10(2):282–290, June 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001881. See errata [304].

VanDerReyden:1954:DAS

[248] D. Van Der Reyden. Design and analysis of soil insecticide field experiments. Biometrics, 10(2):291–297, June 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001882.

Snedecor:1954:Qa

[249] George W. Snedecor. 108. Query. Biometrics, 10(2):298–299, June 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001883.

Snedecor:1954:Qb

[250] George W. Snedecor. 109. Query. Biometrics, 10(2):299–301, June 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001884.

Anonymous:1954:Ab

[251] Anonymous. Abstracts. Biometrics, 10(2):302–314, June 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001885.

Anonymous:1954:FMb

[252] Anonymous. Front matter. Biometrics, 10(2):??, June 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001872.

Grundy:1954:DBT

[253] P. M. Grundy, D. H. Rees, and M. J. R. Healy. Decision between two alternatives — how many experiments? Biometrics, 10(3):317–323, September 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001588.

Yates:1954:AEC

[254] F. Yates. The analysis of experiments containing different crop rotations. Biometrics, 10(3):324–346, September 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001589.

Li:1954:DJD

[255] C. C. Li and Louis Sacks. The derivation of joint distribution and correlation between relatives by the use of stochastic matrices. Biometrics, 10(3): 347–360, September 1954. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001590.

Kimball:1954:SMS

[256] A. W. Kimball and A. S. Householder. A stochastic model for the selection of macronuclear units in paramecium growth. Biometrics, 10(3):361–374, September 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001591.

Bradley:1954:IBR

[257] Ralph Allan Bradley. Incomplete block rank analysis: On the appropriateness of the model for a method of paired comparisons. Biometrics, 10(3): 375–390, September 1954. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001592.

Hopkins:1954:IBR

[258] J. W. Hopkins. Incomplete block rank analysis: Some taste test results. Biometrics, 10(3):391–399, September 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001593.

Nelder:1954:NMP

[259] J. A. Nelder. A note on missing-plot values. Biometrics, 10(3):400–401, September 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001594.

Moore:1954:NTP

[260] P. G. Moore. A note on truncated Poisson distributions. Biometrics, 10 (3):402–406, September 1954. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001595.

Schultz:1954:Q

[261] E. F. Schultz, Jr. 110. Query. Biometrics, 10(3):407–411, September 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001596.

Tukey:1954:Q

[262] John Tukey. 111. Query. Biometrics, 10(3):412–413, September 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001597.

Anonymous:1954:FMc

[263] Anonymous. Front matter. Biometrics, 10(3):??, September 1954. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001587.

Cochran:1954:SMS

[264] William G. Cochran. Some methods for strengthening the common χ^{2} tests. Biometrics, 10(4):417–451, December 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001616.

Kimball:1954:SCF

[265] A. W. Kimball. Short-cut formulas for the exact partition of χ^{2} in contingency tables. Biometrics, 10(4):452–458, December 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001617. See erratum [309].

Almond:1954:NTA

[266] Joyce Almond. A note on the χ^{2} test applied to epidemic chains. Biometrics, 10(4):459–477, December 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001618.

Bross:1954:MT

[267] Irwin Bross. Misclassification in 2 × 2 tables. Biometrics, 10(4):478–486, December 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001619.

Abelson:1954:FPC

[268] Robert M. Abelson and Ralph Allan Bradley. A 2 × 2 factorial with paired comparisons. Biometrics, 10(4):487–502, December 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001620.

Hoblyn:1954:SCD

[269] T. N. Hoblyn, S. C. Pearce, and G. H. Freeman. Some considerations in the design of successive experiments in fruit plantations. Biometrics, 10 (4):503–515, December 1954. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001621.

Graybill:1954:VHR

[270] Franklin Graybill. Variance heterogeneity in a randomized block design. Biometrics, 10(4):516–520, December 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001622.

Hopkins:1954:SOS

[271] J. W. Hopkins. Some observations on sensitivity and repeatability of triad taste difference tests. Biometrics, 10(4):521–530, December 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001623.

Sampford:1954:ERT

[272] M. R. Sampford. The estimation of response-time distribution III truncation and survival. Biometrics, 10(4):531–561, December 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001624.

Moore:1954:Q

[273] Peter G. Moore and John W. Tukey. 112. Query. Biometrics, 10(4): 562–568, December 1954. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001625.

Anonymous:1954:Ac

[274] Anonymous. Abstracts. Biometrics, 10(4):569–570, December 1954. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001626.

Anonymous:1954:BSb

[275] Anonymous. The Biometric Society. Biometrics, 10(4):571–576, December 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001627.

Anonymous:1954:VI

[276] Anonymous. Volume information. Biometrics, 10(4):577–609, December 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001614.

Anonymous:1954:FMd

[277] Anonymous. Front matter. Biometrics, 10(4):??, December 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001615.

Anonymous:1955:FMa

[278] Anonymous. Front matter. Biometrics, 11(1):i–ii, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001477.

Duncan:1955:MRM

[279] David B. Duncan. Multiple range and multiple F tests. Biometrics, 11(1): 1–42, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001478.

Kendall:1955:FCT

[280] M. G. Kendall. Further contributions to the theory of paired comparisons. Biometrics, 11(1):43–62, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001479.

Hopkins:1955:CSP

[281] J. W. Hopkins and N. T. Gridgeman. Comparative sensitivity of pair and triad flavor intensity difference tests. Biometrics, 11(1):63–68, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001480.

Hayman:1955:DGI

[282] B. I. Hayman and K. Mather. The description of genic interactions in continuous variation. Biometrics, 11(1):69–82, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001481.

Holt:1955:QSD

[283] L. B. Holt. Quantitative studies in diphtheria prophylaxis: an attempt to derive a mathematical characterization of the antigenicity of diphtheria prophylactic. Biometrics, 11(1):83–94, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001482.

Robertson:1955:PEQ

[284] Alan Robertson. Prediction equations in quantitative genetics. Biometrics, 11(1):95–98, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001483.

Jessen:1955:DFC

[285] Raymond J. Jessen. Determining the fruit count on a tree by randomized branch sampling. Biometrics, 11(1):99–109, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001484.

Norton:1955:FNM

[286] Horace W. Norton. A further note on missing data. Biometrics, 11(1): 110, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001485.

Tukey:1955:AQ

[287] John W. Tukey. Answer to Query 113. Biometrics, 11(1):111–113, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001486.

Anonymous:1955:Aa

[288] Anonymous. Abstracts. Biometrics, 11(1):114, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001487.

Anonymous:1955:BSa

[289] Anonymous. The Biometric Society. Biometrics, 11(1):115–117, March 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001488.

Anonymous:1955:FMb

[290] Anonymous. Front matter. Biometrics, 11(2):i–ii, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001790.

Schultz:1955:RTD

[291] E. F. Schultz, Jr. Rules of thumb for determining expectations of mean squares in analysis of variance. Biometrics, 11(2):123–135, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001791.

Lowry:1955:VCR

[292] Dorothy C. Lowry. Variance components with reference to genetic population parameters. Biometrics, 11(2):136–148, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001792.

Douglas:1955:FNT

[293] J. B. Douglas. Fitting the Neyman Type A (two parameter) contagious distribution. Biometrics, 11(2):149–173, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001793.

Claringbold:1955:USD

[294] P. J. Claringbold. Use of the simplex design in the study of joint action of related hormones. Biometrics, 11(2):174–185, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001794.

Barraclough:1955:SAM

[295] C. G. Barraclough. Statistical analysis of multiple slope ratio assays. Biometrics, 11(2):186–200, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001795.

Steel:1955:APC

[296] R. G. D. Steel. An analysis of perennial crop data. Biometrics, 11(2):201– 212, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001796.

Bliss:1955:ICB

[297] C. I. Bliss. An introductory course in biometry for graduate students in biology. Biometrics, 11(2):213–224, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001797.

Spicer:1955:TBC

[298] C. C. Spicer. The theory of bacterial constant growth apparatus. Biometrics, 11(2):225–230, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001798.

Huddleston:1955:IMA

[299] Harold F. Huddleston. An inverted matrix approach for determining cropweather regression equations. Biometrics, 11(2):231–236, June 1955. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001799.

Fisher:1955:Q

[300] R. A. Fisher. 114. Query. Biometrics, 11(2):237–238, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001800.

Zelen:1955:Q

[301] Marvin Zelen. 115. Query. Biometrics, 11(2):238–239, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001801.

Federer:1955:Q

[302] W. T. Federer. 116. Query. Biometrics, 11(2):239, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001802.

Anonymous:1955:Ab

[303] Anonymous. Abstracts. Biometrics, 11(2):240–251, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001803.

Anonymous:1955:EUC

[304] Anonymous. Errata: The use of covariance to control gradients in experiments, June, 1954. Biometrics, 11(2):251, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001804. See [247].

Anonymous:1955:FMc

[305] Anonymous. Front matter. Biometrics, 11(3):i–ii, September 1955. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001767.

Hamaker:1955:EDI

[306] H. C. Hamaker. Experimental design in industry. Biometrics, 11(3): 257–286, September 1955. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001768.

Box:1955:EER

[307] G. E. P. Box and P. V. Youle. The exploration and exploitation of response surfaces: an example of the link between the fitted surface and the basic mechanism of the system. Biometrics, 11(3):287–323, September 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001769.

McIntyre:1955:DAT

[308] G. A. McIntyre. Design and analysis of two phase experiments. Biometrics, 11(3):324–334, September 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001770.

Anonymous:1955:ESC

[309] Anonymous. Erratum: Short-cut formulas for the exact partition of χ^{2} in contingency tables. Biometrics, 11(3):334, September 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001771. See [265].

Gridgeman:1955:BTP

[310] N. T. Gridgeman. The Bradley–Terry probability model and preference tasting. Biometrics, 11(3):335–343, September 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001772.

Teissier:1955:DLN

[311] Georges Teissier. Sur la determination de l’Axe d’un nuage rectiligne de points. (French) [On the determination of the axis of a rectilinear cloud of points]. Biometrics, 11(3):344–356, September 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001773.

Reeve:1955:VGC

[312] E. C. R. Reeve. The variance of the genetic correlation coefficient. Biometrics, 11(3):357–374, September 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001774.

Armitage:1955:TLT

[313] P. Armitage. Tests for linear trends in proportions and frequencies. Biometrics, 11(3):375–386, September 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/3001775.

Finney:1955:ETP

[314] D. J. Finney and G. C. Varley. An example of the truncated Poisson distribution. Biometrics, 11(3):387–394, September 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001776.

Anonymous:1955:FMd

[315] Anonymous. Front matter. Biometrics, 11(4):i–ii, December 1955. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001718.

Youden:1955:PRL

[316] W. J. Youden and J. S. Hunter. Partially replicated Latin squares. Biometrics, 11(4):399–405, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001719.

Thompson:1955:RSI

[317] W. A. Thompson, Jr. The relative size of the inter- and intra-block error in an incomplete block design. Biometrics, 11(4):406–426, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001720.

Sprent:1955:NDA

[318] P. Sprent. A note on design and analysis of soil insecticide experiments. Biometrics, 11(4):427–430, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001721.

Outhwaite:1955:CAA

[319] Anne D. Outhwaite and A. Rutherford. Covariance analysis as an alternative to stratification in the control of gradients. Biometrics, 11(4): 431–440, December 1955. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001722.

See erratum [361].

Kramer:1955:AVT

[320] Clyde Young Kramer. On the analysis of variance of a two-way classification with unequal sub-class numbers. Biometrics, 11(4):441–452, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001723.

Ditchburne:1955:MAD

[321] Nell Ditchburne. A method of analysis for a double classification arranged in a triangular table. Biometrics, 11(4):453–464, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001724.

Ipsen:1955:ASB

[322] Johannes Ipsen. Appropriate scores in bio-assays using death-times and survivor symptoms. Biometrics, 11(4):465–480, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001725.

Claringbold:1955:MQA

[323] P. J. Claringbold. Matrices in quantal analysis. Biometrics, 11(4):481–501, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001726.

Finney:1955:LE

[324] D. J. Finney. Letter to the editor. Biometrics, 11(4):502–503, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001727.

Hartley:1955:Q

[325] H. O. Hartley. 118. Query. Biometrics, 11(4):504–505, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001728.

Cochran:1955:CEB

[326] W. G. Cochran. Change in the editorship of Biometrics. Biometrics, 11 (4):506–507, December 1955. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001729.

Anonymous:1955:Ac

[327] Anonymous. Abstracts. Biometrics, 11(4):508–518, December 1955. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001730.

Baily:1955:N

[328] Joshua L. Baily, Jr. Note. Biometrics, 11(4):519, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001731.

Anonymous:1955:NM

[329] Anonymous. News of members. Biometrics, 11(4):519–520, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001732.

Anonymous:1955:I

[330] Anonymous. [Introduction]. Biometrics, 11(4):521–524, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001733.

Anonymous:1955:SP

[331] Anonymous. Scientific program. Biometrics, 11(4):525–526, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001734.

Anonymous:1955:R

[332] Anonymous. Registration. Biometrics, 11(4):527, December 1955. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001735.

Cochran:1955:BPA

[333] William G. Cochran. Brief of presidential address: The 1954 trial of the poliomyelitis vaccine in the United States. Biometrics, 11(4):528–534, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001736.

Anonymous:1955:AP

[334] Anonymous. Abstracts of papers. Biometrics, 11(4):535–555, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001737.

Anonymous:1955:BSb

[335] Anonymous. The Biometric Society. Biometrics, 11(4):556–561, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001738.

Anonymous:1955:VI

[336] Anonymous. Volume information. Biometrics, 11(4):563–574, December 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001717.

Anonymous:1956:FMa

[337] Anonymous. Front matter. Biometrics, 12(1):i–ii, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001570.

Morrison:1956:FRM

[338] Milton Morrison. Fractional replication for mixed series. Biometrics, 12 (1):1–19, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001571. See

correction [378].

DeBaun:1956:BED

[339] Robert M. DeBaun. Block effects in the determination of optimum conditions. Biometrics, 12(1):20–22, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001572.

Truitt:1956:ACC

[340] Jeanne Titus Truitt and H. Fairfield Smith. Adjustment by covariance and consequent tests of significance in split-plot experiments. Biometrics, 12(1):23–39, March 1956. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001573.

Clarke:1956:NCE

[341] Pamela M. Clarke. A note on the combination of estimates of relative potency in multiple assays. Biometrics, 12(1):40–46, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001574.

Watson:1956:MMF

[342] G. S. Watson. Missing and “Mixed-Up” frequencies in contingency tables. Biometrics, 12(1):47–50, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001575.

Ramachandran:1956:CSC

[343] K. V. Ramachandran. Contributions to simultaneous confidence interval estimation. Biometrics, 12(1):51–56, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001576.

Kimura:1956:RGD

[344] Motoo Kimura. Random genetic drift in a tri-allelic locus; exact solution with a continuous model. Biometrics, 12(1):57–66, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001577.

Finney:1956:MAA

[345] D. J. Finney. Multivariate analysis and agricultural experiments. Biometrics, 12(1):67–71, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001578.

Hewlett:1956:RBQ

[346] P. S. Hewlett and R. L. Plackett. The relation between quantal and graded responses to drugs. Biometrics, 12(1):72–78, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001579.

Norton:1956:OLA

[347] H. W. Norton. One likelihood adjustment may be inadequate. Biometrics, 12(1):79–81, March 1956. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001580.

Rigney:1956:Q

[348] J. A. Rigney. 119. Query. Biometrics, 12(1):82–84, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001581.

Lush:1956:Q

[349] J. L. Lush. 120. Query. Biometrics, 12(1):84–88, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001582.

Anonymous:1956:Aa

[350] Anonymous. Abstracts. Biometrics, 12(1):89, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001583.

Anonymous:1956:BSa

[351] Anonymous. The Biometric Society. Biometrics, 12(1):90–92, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001584.

Cavalli-Sforza:1956:VSB

[352] L. L. Cavalli-Sforza. The Varenna Seminar in Biometry. Biometrics, 12 (1):93–96, March 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001585.

Anonymous:1956:FMb

[353] Anonymous. Front matter. Biometrics, 12(2):i–ii, June 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001753.

Graybill:1956:CIV

[354] Franklin A. Graybill, Frank Martin, and George Godfrey. Confidence intervals for variance ratios specifying genetic heritability. Biometrics, 12(2):99–109, June 1956. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001754.

Hartley:1956:PPA

[355] H. O. Hartley. A plan for programming analysis of variance for general purpose computers. Biometrics, 12(2):110–122, June 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001755.

Gurian:1956:NFM

[356] Joan M. Gurian. Note on fitting the multi-hit survival curve. Biometrics, 12(2):123–126, June 1956. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001756.

Kauman:1956:QEN

[357] W. G. Kauman, J. W. Gottstein, and D. Lantican. Quality evaluation by numerical and subjective methods, with application to dried veneer. Biometrics, 12(2):127–153, June 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001757.

Whittle:1956:EAS

[358] P. Whittle. The estimation of age-specific infection rates from a curve of relative infection. Biometrics, 12(2):154–162, June 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001758.

Zippin:1956:ERM

[359] Calvin Zippin. An evaluation of the removal method of estimating animal populations. Biometrics, 12(2):163–189, June 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001759.

Li:1956:CPC

[360] C. C. Li. The concept of path coefficient and its impact on population genetics. Biometrics, 12(2):190–210, June 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001760.

Anonymous:1956:ECA

[361] Anonymous. Erratum: Covariance analysis as an alternative to stratification in the control of gradients. Biometrics, 12(2):210, June 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001761. See [319].

Lewontin:1956:END

[362] R. C. Lewontin and Timothy Prout. Estimation of the number of different classes in a population. Biometrics, 12(2):211–223, June 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001762.

Zyskind:1956:Q

[363] George Zyskind. 121. Query. Biometrics, 12(2):224–225, June 1956. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001763.

Anonymous:1956:Ab

[364] Anonymous. Abstracts. Biometrics, 12(2):226–234, June 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001764.

Anonymous:1956:BSb

[365] Anonymous. The Biometric Society. Biometrics, 12(2):235–237, June 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001765.

Anonymous:1956:FMc

[366] Anonymous. Front matter. Biometrics, 12(3):i–ii, September 1956. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001462.

Thompson:1956:EMP

[367] H. R. Thompson. Extensions to missing plot techniques. Biometrics, 12 (3):241–244, September 1956. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001463.

See corrections [447].

Hodnett:1956:AEA

[368] G. E. Hodnett. The analysis of a 3 × 6 experiment arranged in a quasi-Latin square. Biometrics, 12(3):245–258, September 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001464.

Claringbold:1956:NSF

[369] P. J. Claringbold. A note on the 4^{n} series of factorial experiments. Biometrics, 12(3):259–263, September 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001465.

Rao:1956:SSS

[370] C. Radhakrishna Rao and I. M. Chakravarti. Some small sample tests of significance for a Poisson distribution. Biometrics, 12(3):264–282, September 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/3001466.

Billewicz:1956:MPS

[371] W. Z. Billewicz. Matched pairs in sequential trials for significance of a difference between proportions. Biometrics, 12(3):283–300, September 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001467.

Dykstra:1956:NRA

[372] Otto Dykstra, Jr. A note on the rank analysis of incomplete block designs — applications beyond the scope of existing tables. Biometrics, 12(3): 301–306, September 1956. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001468.

Kramer:1956:EMR

[373] Clyde Young Kramer. Extension of multiple range tests to group means with unequal numbers of replications. Biometrics, 12(3):307–310, September 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/3001469.

Horn:1956:SLE

[374] Henry J. Horn. Simplified LD 50 (or ED 50) calculations. Biometrics, 12 (3):311–322, September 1956. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001470.

Patterson:1956:SMF

[375] H. D. Patterson. A simple method for fitting an asymptotic regression curve. Biometrics, 12(3):323–329, September 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001471.

Pearce:1956:SPE

[376] S. C. Pearce. Some problems of experimental design and technique with perennial crops. Biometrics, 12(3):330–337, September 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001472.

Snedecor:1956:Qa

[377] George W. Snedecor and C. Clark Cockerham. 122. Query. Biometrics, 12 (3):338–339, September 1956. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001473.

Morrison:1956:CFR

[378] Milton Morrison. Correction to “Fractional Replication for Mixed Series”. Biometrics, 12(3):339, September 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001474. See [338].

Anonymous:1956:NAa

[379] Anonymous. News and announcements. Biometrics, 12(3):340, September 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001475.

Anonymous:1956:FMd

[380] Anonymous. Front matter. Biometrics, 12(4):i–ii, December 1956. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001676.

Finney:1956:SAC

[381] D. J. Finney and F. W. Cope. The statistical analysis of a complex experiment involving unintentional constraints. Biometrics, 12(4):345–368, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001677.

Nair:1956:SAS

[382] K. R. Nair. Simplified analysis of singly linked blocks. Biometrics, 12(4): 369–380, December 1956. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001678.

Bliss:1956:RAP

[383] C. I. Bliss, Mary L. Greenwood, and Edna Sakamoto White. A rankit analysis of paired comparisons for measuring the effect of sprays on flavor. Biometrics, 12(4):381–403, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001679.

Horner:1956:TES

[384] Theodore W. Horner and Charles R. Weber. Theoretical and experimental study of self fertilized populations. Biometrics, 12(4):404–414, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001680.

Opsahl:1956:DIL

[385] Birger Opsahl. The discrimination of interactions and linkage in continuous variation. Biometrics, 12(4):415–432, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001681.

Robson:1956:ASG

[386] D. S. Robson. Applications of the k_{4} statistic to genetic variance component analyses. Biometrics, 12(4):433–444, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001682.

Keeping:1956:NWM

[387] E. S. Keeping. Note on Wald’s method of fitting a straight line when both variables are subject to error. Biometrics, 12(4):445–448, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001683.

Masuyama:1956:RAB

[388] M. Masuyama and M. Hatamura. Recent advances in biometry in Japan. Biometrics, 12(4):449–461, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001684.

Hansen:1956:CES

[389] Morris H. Hansen and Joseph Steinberg. Control of errors in surveys. Biometrics, 12(4):462–474, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001685.

Irwin:1956:SPE

[390] J. O. Irwin. The study of the physiological effects of hot climates. Biometrics, 12(4):475–490, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001686.

Bliss:1956:CLM

[391] C. I. Bliss. Confidence limits for measuring the precision of bioassays. Biometrics, 12(4):491–526, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001687.

Snedecor:1956:Qb

[392] George W. Snedecor and G. A. McIntyre. 123. Query. Biometrics, 12(4): 527–532, December 1956. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001688.

Anonymous:1956:Ac

[393] Anonymous. Abstracts. Biometrics, 12(4):533–538, December 1956. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001689.

Anonymous:1956:NAb

[394] Anonymous. News and announcements. Biometrics, 12(4):539–541, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001690.

Anonymous:1956:BSc

[395] Anonymous. The Biometric Society. Biometrics, 12(4):542–544, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001691.

Anonymous:1956:VI

[396] Anonymous. Volume information. Biometrics, 12(4):545–557, December 1956. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001675.

Anonymous:1957:FMa

[397] Anonymous. Front matter. Biometrics, 13(1):i–ii, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001896.

Taylor:1957:IBD

[398] W. B. Taylor. Incomplete block designs with row balance and recovery of inter-block information. Biometrics, 13(1):1–12, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001897.

Kramer:1957:EMR

[399] Clyde Young Kramer. Extension of multiple range tests to group correlated adjusted means. Biometrics, 13(1):13–18, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001898.

Cornish:1957:AKP

[400] E. A. Cornish. An application of the Kronecker product of matrices in multiple regression. Biometrics, 13(1):19–27, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001899.

Berkson:1957:TML

[401] Joseph Berkson. Tables for the maximum likelihood estimate of the logistic function. Biometrics, 13(1):28–34, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001900.

Bliss:1957:BP

[402] C. I. Bliss. Bioassay from a parabola. Biometrics, 13(1):35–50, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001901.

Jackson:1957:ESS

[403] J. Edward Jackson and Mary Fleckenstein. An evaluation of some statistical techniques used in the analysis of paired comparison data. Biometrics, 13(1):51–64, March 1957. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001902.

McGuire:1957:DEC

[404] Judson U. McGuire, Tom A. Brindley, and T. A. Bancroft. The distribution of European corn borer larvae Pyrausta nubilalis (Hbn.), in field corn. Biometrics, 13(1):65–78, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 3001903. See errata and extensions [489].

Chiang:1957:ASP

[405] Chin Long Chiang. An application of stochastic processes to experimental studies on flour beetles. Biometrics, 13(1):79–97, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001904. See corrections [446].

Wallace:1957:UAC

[406] Margaret E. Wallace. The use of affinity in chromosome mapping. Biometrics, 13(1):98–110, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001905.

vanderVeen:1957:HVP

[407] J. H. van der Veen. The heritable variance of F_{1} ×F_{2} progeny means. Biometrics, 13(1):111–112, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001906.

Snedecor:1957:QIT

[408] George W. Snedecor, P. Armitage, and M. J. R. Healy. 124. Query. Interpretation of χ2 tests. Biometrics, 13(1):113–115, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001907.

Smith:1957:NMP

[409] H. Fairfield Smith. 125. Note: Missing plot estimates. Biometrics, 13 (1):115–118, March 1957. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001908.

Anonymous:1957:BSa

[410] Anonymous. The Biometric Society. Biometrics, 13(1):119–122, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001909.

Anonymous:1957:NAa

[411] Anonymous. News and announcements. Biometrics, 13(1):123–125, March 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001910.

Anonymous:1957:FMb

[412] Anonymous. Front matter. Biometrics, 13(2):i–ii, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527795.

Pearce:1957:RBS

[413] S. C. Pearce and D. A. Holland. Randomized branch sampling for estimating fruit number. Biometrics, 13(2):127–130, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527796.

Sheps:1957:ERP

[414] Mindel C. Sheps and Paul L. Munson. The error of replicated potency estimates in a biological assay method of the parallel line type. Biometrics, 13(2):131–148, June 1957. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527797.

Meynell:1957:ILP

[415] G. G. Meynell. Inherently low precision of infectivity titrations using a quantal response. Biometrics, 13(2):149–163, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527798.

Duncan:1957:MRT

[416] David B. Duncan. Multiple range tests for correlated and heteroscedastic means. Biometrics, 13(2):164–176, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527799.

Ipsen:1957:ASR

[417] Johannes Ipsen, Jr. Appropriate scores for reaction categories dependent on two variables. Biometrics, 13(2):177–182, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527800.

Blackith:1957:PSA

[418] R. E. Blackith. Polymorphism in some Australian locusts and grasshoppers. Biometrics, 13(2):183–196, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527801.

Kramer:1957:EIB

[419] Clyde Young Kramer and Ralph Allan Bradley. Examples of intra-block analysis for factorials in group divisible, partially balanced, incomplete block designs. Biometrics, 13(2):197–224, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527802.

Li:1957:RLR

[420] C. C. Li. Repeated linear regression and variance components of a population with binomial frequencies. Biometrics, 13(2):225–233, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527803.

Cornish:1957:ACB

[421] E. A. Cornish. Appreciation: Chester I. Bliss. Biometrics, 13(2):234, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527804.

Kempthorne:1957:QAP

[422] Oscar Kempthorne. 126. Query: Arrangements of pots in greenhouse experiments. Biometrics, 13(2):235–237, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527805.

Anonymous:1957:Aa

[423] Anonymous. Abstracts. Biometrics, 13(2):238–246, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527806.

Anonymous:1957:BSb

[424] Anonymous. The Biometric Society. Biometrics, 13(2):247–256, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527807.

Anonymous:1957:NAb

[425] Anonymous. News and announcements. Biometrics, 13(2):257–258, June 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527808.

Anonymous:1957:FMc

[426] Anonymous. Front matter. Biometrics, 13(3):i–ii, September 1957. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527915.

Cochran:1957:ACN

[427] William G. Cochran. Analysis of covariance: Its nature and uses. Biometrics, 13(3):261–281, September 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527916.

Smith:1957:IAT

[428] H. Fairfield Smith. Interpretation of adjusted treatment means and regressions in analysis of covariance. Biometrics, 13(3):282–308, September 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527917.

Zelen:1957:ACI

[429] Marvin Zelen. The analysis of covariance for incomplete block designs. Biometrics, 13(3):309–332, September 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527918.

Federer:1957:VCA

[430] Walter T. Federer. Variance and covariance analyses for unbalanced classifications. Biometrics, 13(3):333–362, September 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527919.

Wilkinson:1957:ACI

[431] G. N. Wilkinson. The analysis of covariance with incomplete data. Biometrics, 13(3):363–372, September 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527920.

Finney:1957:SBC

[432] D. J. Finney. Stratification, balance, and covariance. Biometrics, 13(3): 373–386, September 1957. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527921.

Coons:1957:ACM

[433] Irma Coons. The analysis of covariance as a missing plot technique. Biometrics, 13(3):387–405, September 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527922.

Anonymous:1957:Ab

[434] Anonymous. Abstracts. Biometrics, 13(3):406–422, September 1957. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527923.

Anonymous:1957:BSc

[435] Anonymous. The Biometric Society. Biometrics, 13(3):423–429, September 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527924.

Anonymous:1957:NAc

[436] Anonymous. News and announcements. Biometrics, 13(3):430, September 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527925.

Anonymous:1957:FMd

[437] Anonymous. Front matter. Biometrics, 13(4):i–ii, December 1957. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527969.

Maritz:1957:MDS

[438] J. S. Maritz. A model for the deterioration in strength of materials due to fungal attack. Biometrics, 13(4):433–441, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527970.

Robertson:1957:OGS

[439] Alan Robertson. Optimum group size in progeny testing and family selection. Biometrics, 13(4):442–450, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527971.

Corsten:1957:PEV

[440] L. C. A. Corsten. Partition of experimental vectors connected with multinomial distributions. Biometrics, 13(4):451–484, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527972.

Turner:1957:DRB

[441] Malcolm E. Turner. The distribution of red blood cells in the hemacytometer. Biometrics, 13(4):485–495, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527973.

Bradley:1957:CSS

[442] R. A. Bradley and D. E. W. Schumann. The comparison of the sensitivities of similar experiments: Applications. Biometrics, 13(4):496–510, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527974.

Harter:1957:ERS

[443] H. Leon Harter. Error rates and sample sizes for range tests in multiple comparisons. Biometrics, 13(4):511–536, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527975.

Madow:1957:NSS

[444] William G. Madow. 127 Note: Some simple methods of computing parameters in the analysis of variance. Biometrics, 13(4):537–540, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527976.

Matern:1957:NRC

[445] Bertil Matérn. 128 Note: a routine for computing the degrees of freedom in analysis of variance. Biometrics, 13(4):541–543, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527977.

Anonymous:1957:CPA

[446] Anonymous. Corrections to papers: An application of stochastic processes to experimental studies on flour beetles. Biometrics, 13(4):543, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527978. See [405].

Anonymous:1957:CPE

[447] Anonymous. Corrections to papers: Extensions to missing plot techniques. Biometrics, 13(4):543, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527980. See [367].

Anonymous:1957:CPF

[448] Anonymous. Corrections to papers: Fixed-sample-size analysis of sequential observations. Biometrics, 13(4):543, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527979. See [229].

Cornish:1957:A

[449] E. A. Cornish. Appreciation. Biometrics, 13(4):544, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527981.

Anonymous:1957:Ac

[450] Anonymous. Abstracts. Biometrics, 13(4):545–552, December 1957. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527982.

Anonymous:1957:BSd

[451] Anonymous. The Biometric Society. Biometrics, 13(4):553–556, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527983.

Anonymous:1957:NAd

[452] Anonymous. News and announcements. Biometrics, 13(4):557–558, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527984.

Anonymous:1957:VI

[453] Anonymous. Volume information. Biometrics, 13(4):559–570, December 1957. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527968.

Anonymous:1958:FMa

[454] Anonymous. Front matter. Biometrics, 14(1):i–ii, March 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527725.

Rao:1958:SSM

[455] C. Radhakrishna Rao. Some statistical methods for comparison of growth curves. Biometrics, 14(1):1–17, March 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527726.

Bross:1958:HUR

[456] Irwin D. J. Bross. How to use ridit analysis. Biometrics, 14(1):18–38, March 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527727.

Ferris:1958:VMC

[457] George E. Ferris. The k-visit method of consumer testing. Biometrics, 14(1):39–49, March 1958. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527728.

Fitzpatrick:1958:DID

[458] Robert Fitzpatrick. The detection of individual differences in accident susceptibility. Biometrics, 14(1):50–68, March 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527729.

Kelleher:1958:PEV

[459] Therese Kelleher, H. F. Robinson, and R. E. Comstock. Precision of estimates of variance components. Biometrics, 14(1):69–77, March 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527730.

Sukhatme:1958:STE

[460] P. V. Sukhatme. Sampling technique for estimating the catch of sea fish in India. Biometrics, 14(1):78–96, March 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527731.

Sprott:1958:MML

[461] D. A. Sprott. The method of maximum likelihood applied to the Poisson binomial distribution. Biometrics, 14(1):97–106, March 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527732.

Smith:1958:MAC

[462] H. Fairfield Smith. A multivariate analysis of covariance. Biometrics, 14 (1):107–127, March 1958. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527733.

Stuart:1958:NIS

[463] Alan Stuart. 129. Note: Iterative solutions of likelihood equations. Biometrics, 14(1):128–130, March 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527734.

Anonymous:1958:Aa

[464] Anonymous. Abstracts. Biometrics, 14(1):131–146, March 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527735.

Anonymous:1958:BSa

[465] Anonymous. The Biometric Society. Biometrics, 14(1):147–152, March 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527736.

Anonymous:1958:NAa

[466] Anonymous. News and announcements. Biometrics, 14(1):153–157, March 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527737.

Anonymous:1958:FMb

[467] Anonymous. Front matter. Biometrics, 14(2):i–ii, June 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527781.

Simpson:1958:ESM

[468] H. R. Simpson. The effect of sterilised males on a natural tsetse fly population. Biometrics, 14(2):159–173, June 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527782.

Hartley:1958:MLE

[469] H. O. Hartley. Maximum likelihood estimation from incomplete data. Biometrics, 14(2):174–194, June 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 2527783. See corrections [505].

Harris:1958:PSB

[470] Eugene K. Harris. On the probability of survival of bacteria in sea water. Biometrics, 14(2):195–206, June 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527784.

Hatheway:1958:EER

[471] W. H. Hatheway and E. J. Williams. Efficient estimation of the relationship between plot size and the variability of crop yields. Biometrics, 14(2):207–222, June 1958. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527785.

Kastenbaum:1958:ERF

[472] Marvin A. Kastenbaum. Estimation of relative frequencies of four sperm types in Drosophila melanogaster. Biometrics, 14(2):223–228, June 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527786.

Gurland:1958:GCC

[473] John Gurland. A generalized class of contagious distributions. Biometrics, 14(2):229–249, June 1958. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527787.

Vickery:1958:REA

[474] R. E. Vickery. Recent experiences with area sampling for agricultural statistics. Biometrics, 14(2):250–256, June 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527788.

Wilkinson:1958:EMV

[475] G. N. Wilkinson. Estimation of missing values for the analysis of incomplete data. Biometrics, 14(2):257–286, June 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527789.

Grandage:1958:QOC

[476] A. Grandage. 130. Query: Orthogonal coefficients for unequal intervals. Biometrics, 14(2):287–289, June 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527790.

Anonymous:1958:Ab

[477] Anonymous. Abstracts. Biometrics, 14(2):290–294, June 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527791.

Anonymous:1958:BSb

[478] Anonymous. The Biometric Society. Biometrics, 14(2):295–304, June 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527792.

Anonymous:1958:NAb

[479] Anonymous. News and announcements. Biometrics, 14(2):305–307, June 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527793.

Anonymous:1958:FMc

[480] Anonymous. Front matter. Biometrics, 14(3):i–ii, September 1958. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527876.

Bliss:1958:FDB

[481] C. I. Bliss. The first decade of the biometric society. Biometrics, 14(3): 309–329, September 1958. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527877.

Meier:1958:RMS

[482] Paul Meier, Spencer M. Free, Jr., and George L. Jackson. Reconsideration of methodology in studies of pain relief. Biometrics, 14(3):330–342, September 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527878.

Mather:1958:IGE

[483] Kenneth Mather and R. Morley Jones. Interaction of genotype and environment in continuous variation: I. Description. Biometrics, 14(3):343– 359, September 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527879.

Wilkinson:1958:AVD

[484] G. N. Wilkinson. The analysis of variance and derivation of standard errors for incomplete data. Biometrics, 14(3):360–384, September 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527880.

Skellam:1958:MFU

[485] J. G. Skellam. The mathematical foundations underlying the use of line transects in animal ecology. Biometrics, 14(3):385–400, September 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527881.

Fleischer:1958:MEA

[486] Jack Fleischer, Daniel G. Horvitz, J. Malcolm Airth, and A. L. Finkner. Measurement errors associated with obtaining acreage estimates of cotton fields. Biometrics, 14(3):401–407, September 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527882.

Bechhofer:1958:SMD

[487] Robert E. Bechhofer. A sequential multiple-decision procedure for selecting the best one of several normal populations with a common unknown variance, and its use with various experimental designs. Biometrics, 14(3): 408–429, September 1958. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527883.

Daniel:1958:NVO

[488] Cuthbert Daniel. 131 Note: On varying one factor at a time. Biometrics, 14(3):430–431, September 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527884.

McGuire:1958:NEE

[489] Judson U. McGuire, Tom A. Brindley, and T. A. Bancroft. 132 Note: Errata and extensions for “The distribution of the European corn borer larvae Pyrausta nubilalis (Hbn.), in field corn”. Biometrics, 14(3):432–434, September 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527885. See [404].

Anonymous:1958:Ac

[490] Anonymous. Abstracts. Biometrics, 14(3):435–441, September 1958. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527886.

Anonymous:1958:BSc

[491] Anonymous. The Biometric Society. Biometrics, 14(3):442–447, September 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527887.

Anonymous:1958:NAc

[492] Anonymous. News and announcements. Biometrics, 14(3):448–450, September 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527888.

Spehl:1958:ORL

[493] P. Spehl. Obituary: Roger Laurent. Biometrics, 14(3):451, September 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527889.

Anonymous:1958:FMd

[494] Anonymous. Front matter. Biometrics, 14(4):i–ii, December 1958. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527512.

Hodges:1958:FLM

[495] J. L. Hodges, Jr. Fitting the logistic by maximum likelihood. Biometrics, 14(4):453–461, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527513.

White:1958:MGR

[496] Robert F. White and Joseph G. Graca. Multinomially grouped response times for the quantal response bioassay. Biometrics, 14(4):462–488, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527514.

Jones:1958:IGE

[497] R. Morley Jones and Kenneth Mather. Interaction of genotype and environment in continuous variation: II. Analysis. Biometrics, 14(4):489–498, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527515.

Cox:1958:ETT

[498] C. P. Cox. Experiments with two treatments per experimental unit in the presence of an individual covariate. Biometrics, 14(4):499–512, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527516.

Lejeune:1958:SPM

[499] Jérôme Lejeune. Sur une solution “A priori” de la methode “A posteriori” de Haldane. (French) [On an “A priori” solution of the “A posteriori” method of Haldane]. Biometrics, 14(4):513–520, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527517.

Gomes:1958:JAE

[500] Frederico Pimentel Gomes and Rubens Foot Guimarães. Joint analysis of experiments in complete randomised blocks with some common treatments. Biometrics, 14(4):521–526, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527518.

Alling:1958:AHP

[501] David W. Alling. The after-history of pulmonary tuberculosis: a stochastic model. Biometrics, 14(4):527–547, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527519.

Gridgeman:1958:AQR

[502] N. T. Gridgeman. Application of quantal-response theory to the crosscomparison of taste-stimuli intensities. Biometrics, 14(4):548–557, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527520.

Hamaker:1958:NHC

[503] H. C. Hamaker. 133. Note: On hemacytometer counts. Biometrics, 14(4): 558–559, December 1958. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527521.

Anonymous:1958:QCS

[504] Anonymous. 134. Query: Chi-squares of Bartlett, Mood, and Lancaster in a 2^{3} contingency table. Biometrics, 14(4):560–562, December 1958.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527522.

Anonymous:1958:CPM

[505] Anonymous. Corrections to papers: Maximum likelihood estimation from incomplete data. Biometrics, 14(4):562, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527523. See [469].

Anonymous:1958:Ad

[506] Anonymous. Abstracts. Biometrics, 14(4):563–574, December 1958. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527524.

Anonymous:1958:BSd

[507] Anonymous. The Biometric Society. Biometrics, 14(4):575–582, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527525.

Anonymous:1958:NAd

[508] Anonymous. News and announcements. Biometrics, 14(4):583–585, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527526.

Anonymous:1958:VI

[509] Anonymous. Volume information. Biometrics, 14(4):587–598, December 1958. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527511.

Armitage:1959:HVD

[510] Peter Armitage. Host variability in dilution experiments. Biometrics, 15 (1):1–9, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527597.

Kempthorne:1959:RSI

[511] Oscar Kempthorne and Arne W. Nordskog. Restricted selection indices. Biometrics, 15(1):10–19, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527598.

Parsons:1959:EAT

[512] P. A. Parsons. Equilibria in auto-tetraploids under natural selection for a simplified model of viabilities. Biometrics, 15(1):20–29, March 1959.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527599.

Patterson:1959:ANR

[513] H. D. Patterson. The analysis of a non-replicated experiment involving a single four-course rotation of crops. Biometrics, 15(1):30–59, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527600.

Curnow:1959:ATP

[514] R. N. Curnow. The analysis of a two phase experiment. Biometrics, 15(1):60–73, March 1959. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527601.

Folks:1959:AQR

[515] John Leroy Folks. Analysis of quadruple rectangular lattice designs. Biometrics, 15(1):74–86, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527602.

Sheps:1959:ESM

[516] Mindel C. Sheps. An examination of some methods of comparing several rates or proportions. Biometrics, 15(1):87–97, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527603.

Leech:1959:AEG

[517] F. B. Leech and M. J. R. Healy. The analysis of experiments on growth rate. Biometrics, 15(1):98–106, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor. org/stable/2527604. See corrections [570].

Kastenbaum:1959:CCS

[518] Marvin A. Kastenbaum and Donald E. Lamphiear. Calculation of chisquare to test the no three-factor interaction hypothesis. Biometrics, 15 (1):107–115, March 1959. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527605.

Patterson:1959:EPC

[519] H. D. Patterson and H. L. Lucas. Extra-period change-over designs. Biometrics, 15(1):116–132, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527606.

Box:1959:QRN

[520] G. E. P. Box. 135. Query: Replication of non-center points in the rotatable and near-rotatable central composite design. Biometrics, 15(1): 133–135, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527607.

Williams:1959:QSD

[521] E. J. Williams. 136. Query: Significance of difference between two nonindependent correlation coefficients. Biometrics, 15(1):135–136, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527608.

Anonymous:1959:Aa

[522] Anonymous. Abstracts. Biometrics, 15(1):137–167, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527609.

Anonymous:1959:BSa

[523] Anonymous. The Biometric Society. Biometrics, 15(1):168–179, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527610.

Anonymous:1959:NAa

[524] Anonymous. News and announcements. Biometrics, 15(1):180–185, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527611.

Anonymous:1959:FMa

[525] Anonymous. Front matter. Biometrics, 15(1):??, March 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527596.

Robson:1959:SMC

[526] D. S. Robson. A simple method for constructing orthogonal polynomials when the independent variable is unequally spaced. Biometrics, 15(2): 187–191, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527668.

Henderson:1959:EEG

[527] C. R. Henderson, Oscar Kempthorne, S. R. Searle, and C. M. von Krosigk. The estimation of environmental and genetic trends from records subject to culling. Biometrics, 15(2):192–218, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527669.

Robertson:1959:EDE

[528] Alan Robertson. Experimental design in the evaluation of genetic parameters. Biometrics, 15(2):219–226, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527670.

Schwartz:1959:DFA

[529] Lorraine Schwartz and Stanley Wearden. A distribution-free asymptotic method of estimating, testing, and setting confidence limits for heritability. Biometrics, 15(2):227–235, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527671.

Turner:1959:RAC

[530] Malcolm E. Turner and Charles D. Stevens. The regression analysis of causal paths. Biometrics, 15(2):236–258, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527672.

Roy:1959:CTR

[531] J. Roy. A class of two replicate incomplete block designs. Biometrics, 15(2):259–269, June 1959. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527673.

Milne:1959:CSA

[532] A. Milne. The centric systematic area-sample treated as a random sample. Biometrics, 15(2):270–297, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527674.

Gridgeman:1959:SIS

[533] N. T. Gridgeman. Sensory item sorting. Biometrics, 15(2):298–306, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527675.

Claringbold:1959:OCS

[534] P. J. Claringbold. Orthogonal contrasts in slope ratio investigations. Biometrics, 15(2):307–322, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527676.

Kastenbaum:1959:NCI

[535] Marvin A. Kastenbaum. 137. Note: a confidence interval on the abscissa of the point of intersection of two fitted linear regressions. Biometrics, 15(2):323–324, June 1959. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527677.

Irwin:1959:NEM

[536] J. O. Irwin. 138. Note: On the estimation of the mean of a Poisson distribution from a sample with the zero class missing. Biometrics, 15(2): 324–326, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527678.

Webster:1959:QDR

[537] John T. Webster. 139. Query: Differential regression. Biometrics, 15(2): 326–329, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2527679.

Anonymous:1959:Ab

[538] Anonymous. Abstracts. Biometrics, 15(2):330–341, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527680.

Anonymous:1959:BSb

[539] Anonymous. The Biometric Society. Biometrics, 15(2):342–358, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527681.

Anonymous:1959:NAb

[540] Anonymous. New and announcements. Biometrics, 15(2):359–360, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527682.

Anonymous:1959:FMb

[541] Anonymous. Front matter. Biometrics, 15(2):??, June 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527667.

Irwin:1959:BMP

[542] J. O. Irwin. Biometric method, past, present, and future. Biometrics, 15 (3):363–375, September 1959. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527740.

Rendel:1959:OGS

[543] J. M. Rendel. Optimum group size in half-sib family selection. Biometrics, 15(3):376–381, September 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527741.

Gridgeman:1959:PCT

[544] N. T. Gridgeman. Pair comparison, with and without ties. Biometrics, 15 (3):382–388, September 1959. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527742.

Geisser:1959:MTT

[545] Seymour Geisser. A method for testing treatment effects in the presence of learning. Biometrics, 15(3):389–395, September 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527743.

Chassan:1959:DCS

[546] J. B. Chassan. On the development of clinical statistical systems for psychiatry. Biometrics, 15(3):396–404, September 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527744.

Schumann:1959:CSS

[547] D. E. W. Schumann and R. A. Bradley. The comparison of the sensitivities of similar experiments: Model II of the analysis of variance. Biometrics, 15 (3):405–416, September 1959. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2527745.

See corrections [571].

Wearden:1959:UPF

[548] Stanley Wearden. The use of the power function to determine an adequate number of progeny per sire in a genetic experiment involving half-sibs. Biometrics, 15(3):417–423, September 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527746.

Harris:1959:CLL

[549] Eugene K. Harris. Confidence limits for the LD 50 using the moving average-angle method. Biometrics, 15(3):424–432, September 1959. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527747.

Gjeddebaek:1959:CSG

[550] N. F. Gjeddebaek. Contribution to the study of grouped observations. IV. Some comments on simple estimates. Biometrics, 15(3):433–439, September 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527748.

Watson:1959:SRR

[551] G. S. Watson. Some recent results in chi-square goodness-of-fit tests. Biometrics, 15(3):440–468, September 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527749.

Robertson:1959:SVG

[552] Alan Robertson. The sampling variance of the genetic correlation coefficient. Biometrics, 15(3):469–485, September 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527750.

Finney:1959:QTD

[553] D. J. Finney and David Hogben. 140. Query: Test of difference between treatment and control with multiple replications of control and a missing plot. Biometrics, 15(3):486–487, September 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527751.

Anonymous:1959:Ac

[554] Anonymous. Abstracts. Biometrics, 15(3):489–497, September 1959. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527752.

Anonymous:1959:BSc

[555] Anonymous. The Biometric Society. Biometrics, 15(3):498–505, September 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2527753.

Anonymous:1959:NAc

[556] Anonymous. News and announcements. Biometrics, 15(3):506–510, September 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527754.

Anonymous:1959:FMc

[557] Anonymous. Front matter. Biometrics, 15(3):??, September 1959. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527739.

Nordskog:1959:NOG

[558] A. W. Nordskog. Note on optimum group size for progeny tests. Biometrics, 15(4):513–517, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527649.

Mode:1959:PGV

[559] C. J. Mode and H. F. Robinson. Pleiotropism and the genetic variance and covariance. Biometrics, 15(4):518–537, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527650.

Bodmer:1959:SED

[560] W. F. Bodmer. A significantly extreme deviate in data with a nonsignificant heterogeneity chi square. Biometrics, 15(4):538–542, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527651.

Graybill:1959:CUE

[561] Franklin A. Graybill and R. B. Deal. Combining unbiased estimators. Biometrics, 15(4):543–550, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527652.

Guttman:1959:IUS

[562] Louis Guttman and Ruth Guttman. An illustration of the use of stochastic approximation. Biometrics, 15(4):551–559, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527653.

Steel:1959:MCR

[563] Robert G. D. Steel. A multiple comparison rank sum test: Treatments versus control. Biometrics, 15(4):560–572, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527654.

Ashford:1959:AAD

[564] J. R. Ashford. An approach to the analysis of data for semi-quantal responses in biological assay. Biometrics, 15(4):573–581, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527655.

Anderson:1959:UCT

[565] R. L. Anderson. Use of contingency tables in the analysis of consumer preference studies. Biometrics, 15(4):582–590, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527656.

Hewlett:1959:UTQ

[566] P. S. Hewlett and R. L. Plackett. A unified theory for quantal responses to mixtures of drugs: Non-interactive action. Biometrics, 15(4):591–610, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527657.

Hartley:1959:SCD

[567] H. O. Hartley. Smallest composite designs for quadratic response surfaces. Biometrics, 15(4):611–624, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2527658.

Kupperman:1959:NRS

[568] Morton Kupperman. 141. Note: a rapid significance test for contingency tables. Biometrics, 15(4):625–628, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527659.

Jebe:1959:QQR

[569] Emil H. Jebe. 142. Query: On a quail roadside count technique. Biometrics, 15(4):628–631, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527660.

Anonymous:1959:CAE

[570] Anonymous. Corrections: The analysis of experiments on growth rate. Biometrics, 15(4):631, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2527661. See [517].

Anonymous:1959:CCS

[571] Anonymous. Corrections: The comparison of the sensitivities of similar experiments: Model II of the analysis of variance. Biometrics, 15(4):631, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527662. See [547].

Anonymous:1959:Ad

[572] Anonymous. Abstracts. Biometrics, 15(4):632–642, December 1959. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527663.

Tukey:1959:LPM

[573] John W. Tukey. Little pieces of mixed factorials (abstract 627). Biometrics, 15(4):641–642, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2527663.

Anonymous:1959:BSd

[574] Anonymous. The Biometric Society. Biometrics, 15(4):643–647, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527664.

Anonymous:1959:NAd

[575] Anonymous. News and announcements. Biometrics, 15(4):648–652, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2527665.

Anonymous:1959:VI

[576] Anonymous. Volume information. Biometrics, 15(4):653–670, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2527647.

Anonymous:1959:FMd

[577] Anonymous. Front matter. Biometrics, 15(4):??, December 1959. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2527648.

Sandvik:1978:EIM

[578] Leiv Sandvik. Erratum: “Inverse, multiple and sequential sample censuses” (Biometrics 8 (1952), 286–306) by D. G. Chapman. Biometrics, 34(3):523, September 1978. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [156].